### v1

# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

Abstract
In this work we solve, for given bounded operators $B,C$ and Hilbert–Schmidt operator $M$ acting on potentially infinite-dimensional separable Hilbert spaces, the reduced rank approximation problem,
$\min\{\lVert M-BXC\rVert_{HS}:\ \dim\operatorname{ran}{X}\leq r\}.$
This extends the result of Sondermann (Statistische Hefte, 1986) and Friedland and Torokhti (SIAM J. Matrix Analysis and Applications, 2007), which studies this problem in the case of matrices $M$, $B$, $C$, $X$, and the analysis involves the Moore–Penrose inverse. In classical approximation problems which can be solved by the singular value decomposition or Moore–Penrose inverse, the solution satisfies a minimal norm property. Friedland and Torokhti also state such a minimal norm property of the solution. However, we show that this minimal norm property does not hold in general and give a modified minimality property that is satisfied in general. We also show that the solution may be discontinuous in infinite-dimensional settings. Necessary and sufficient conditions for continuity of the solutions are given and continuous approximations are constructed when such conditions are not met. Finally, we show that our results can be used in signal processing, reduced rank regression and linear operator learning under a rank constraint.
Keywords: rank-constrained approximation, Hilbert-Schmidt operators, infinite-dimensional analysis, unbounded operators, signal processing, reduced rank regression, linear operator learning
1 Introduction
We study the following reduced rank approximation in Hilbert–Schmidt norm $\lVert\cdot\rVert_{HS}$,
$\min\{\lVert M-BXC\rVert_{HS}:\ \dim\operatorname{ran}{X}\leq r\},$ (1)
where $r\in\mathbb{N}$ and $M$, $B$, $X$, $C$ are suitable linear operators. That is, we want to find the best approximation to $M$ of the form $BXC$ for given $B$ and $C$ in Hilbert–Schmidt norm. Such approximations have also been called ‘generalised approximations’, see e.g. [20, 34]. In the finite-dimensional case, the operators can be represented as matrices by expressing them in orthonormal bases. The Hilbert–Schmidt norm then becomes the Frobenius norm in the matrix formulation of the problem.
In the finite-dimensional matrix case, the generalised approximation problem is well studied, cf. [38, 20, 40]. In [20, Theorem 2.1] and [38] a solution of the finite-dimensional matrix problem is given. Furthermore, conditions for uniqueness of the solution are given. A minimality property of the solution is also given in [20, Theorem 2.1], namely that it should have minimal norm among all solutions of the approximation problem. In [40] an error analysis of the solution proposed by [20, Theorem 2.1] is given.
Next, we state the main contributions of this work, review some of the literature on applications of generalised rank-constrained approximations — with or without the minimal norm property — introduce notation and outline the rest of the paper.
1.1 Contributions of this work
The first aim of this work is to show that a conclusion in [20, Theorem 2.1] — namely, that the solution of a certain generalised approximation problem satisfies a minimal norm property — is not in general correct. We therefore propose a refinement of [20, Theorem 2.1] by modifying the minimal norm property. This modified property is always satisfied by the solution given in [20, Theorem 2.1] and can still be interpreted as a minimality property.
The second aim of this work is to generalise the refined statement to an infinite-dimensional setting. We show that the generalised approximation problem does on the one hand not always admit an optimal solution. On the other hand, even when solutions exist, they are no longer bounded in general, due to the Moore–Penrose inverse of $C$ that appears in the solution expression. It is even possible that some solutions are continuous, i.e. bounded, and have finite Hilbert–Schmidt norm, while others are discontinuous, i.e. unbounded, and as such do not even have finite operator norm. Formulation and analysis of the problem thus requires the use of unbounded operators. While the Moore–Penrose inverse is defined also for certain kinds of unbounded operators such as closed range operators, we analyse the problem assuming that $B$ and $C$ do not only have closed range but are bounded linear operators. Two natural questions arise: when does a solution exist, and when are the solutions bounded? We give necessary and sufficient conditions for a solution to the generalised approximation problem to exist in infinite dimensions. When it does not exist, i.e. when the infimum is not attained, we construct a sequence of approximate solutions that converges to this infimum. We shall also give necessary and sufficient conditions for all solutions to be bounded. When no bounded solutions exist, we propose a procedure to construct bounded approximations.
The third aim of this work is to demonstrate how our results can be used in relevant applications of current research interest such as operator learning [30] and in particular linear operator learning [17, 36], reduced rank regression [27, 28, 62] and signal processing [51, 54]. In such applications, there is a direct interest in an infinite-dimensional formulation. For example, as mentioned in [36], the setting of infinite-dimensional linear operator learning has direct applications to functional regression with functional response and conditional mean embeddings. In such an operator learning setting, given samples from the distribution of Hilbert space-valued random variables $x$ and $y$, the goal is to learn a linear operator $A$ that satisfies
$\min\{\mathbb{E}\lVert y-Ax\rVert^{2},A\in\mathcal{A}\},$ (2)
where $\mathcal{A}$ can equal for example the bounded linear or Hilbert–Schmidt operators. This is also studied as a linear signal processing problem in [51], where $x$ and $y$ are $\mathbb{R}^{m}$ or $\mathbb{C}^{m}$ valued. In [51] $x$, $y$, and $A$ are also allowed to be reweighted and the rank of $A$ can be restricted. Under such a rank restriction, the above problem can in fact also be seen as a reduced rank regression problem similar to [27, 62]. We show that these problems can be rewritten in formulation (1). In [28], a reduced rank regression problem of the form (1) is considered where the Hilbert–Schmidt norm is replaced by the operator norm. Applications from the perspective of linear operator learning, signal processing and reduced rank regression thus all motivate the study of a problem of form (2). Using the obtained solution of problem (1) in the infinite-dimensional setting, we therefore analyse problem (2), where $\mathcal{A}$ is the space of finite-rank operators between separable Hilbert spaces. We also allow for reweighting of $x,y$ and $A$ as in [51], and give necessary and sufficient conditions for a solution to exist, constructing also an explicit solution that satisfies a minimality property. In this application, we show that our proposed minimality property of the solutions corresponds to a maximal kernel property that naturally occurs in linear operator learning.
Even in finite-dimensional settings, the study of an infinite-dimensional analogue can be beneficial. Infinite-dimensional problems are often discretized and transformed to finite-dimensional generalised rank-constrained approximation problems. This happens for example in the study of inverse problems governed by partial differential equations, e.g. [12]. An infinite-dimensional version of the generalised approximation problem would then enable the analysis of such problems directly in their infinite-dimensional formulation, enabling us to draw conclusions that are independent of the discretization dimension.
1.2 Related literature
In finite-dimensional settings, the reduced rank approximation problem (1) has been applied in the fields of signal processing [51, 52, 55, 60, 44, 54, 45, 58, 56, 46, 42, 50, 44, 43, 21, 57, 59], data transmission systems [61, 41], prescribed eigenvalue approximation [34], regularized inverse problems [12], subspace estimation in computer vision [33], neuroimaging [32], classification [35], recommender systems [66, 9], image processing [10] and in distributed memory computing [7]. It is also used for tensor approximation [37, 3], low-rank posterior approximation for linear-Gaussian inverse problems [49, 48, 6], randomised matrix approximation algorithms [39, 2], regression [63], Bayes-risk minimisation in linear inverse problems [11], generative modeling [31] and fast least squares problems in statistical learning [64]. The analogue of (1) without the rank constraint is also used in other numerical linear algebra tasks [22]. Further applications are mentioned in [8].
For treatment of a similar problem to (1) in finite dimensions, in which the operator norm is used instead of the Hilbert–Schmidt norm, see [47, 28].
In [20, Theorem 2.1] a solution is proposed for the reduced rank approximation problem, and a minimal norm property is stated for this solution. This minimal norm property is reiterated in other works, [7, 53, 34, 54, 45, 12, 19], and used in [23, 65, 51, 61, 41, 59, 33]. Note that [40], which adds an error analysis, does not state the minimal norm property. The minimal norm property also has direct use in applications, as for example in the context of dynamic scene reconstruction from image measurements of [33]. There, it is mentioned that the choice of solutions has a large effect on the quality of the reconstruction, and a minimal norm solution favours points close to the camera plane.
Imposing a minimality property has the consequence of mitigating nonuniqueness of the solution. This can also be achieved by imposing a regularizing term $\alpha^{2}\lVert X\rVert_{HS}^{2}$. This is the approach of [13] for a similar reduced rank approximation problem in the context of inverse problems.
1.3 Notation
For a matrix $Y\in\mathbb{C}^{s\times t}$, we denote by $(Y)_{r}$ a rank-$r$ truncated singular value decomposition (SVD) of $Y$. Thus if $Y=U\Sigma V^{*}$ is an SVD of $Y$, then $(Y)_{r}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{*}$ where $\sigma_{i}=\Sigma_{ii}$ are nonincreasing and $u_{i}$ and $v_{i}$ are the $i$-th column of $U$ and $V$ respectively. The Frobenius norm of $Y$ is denoted as $\lVert Y\rVert_{F}$. By $Y^{\dagger}\in\mathbb{C}^{t\times s}$ we denote the Moore–Penrose inverse, also known as the pseudo-inverse or generalised inverse. We define $P_{Y,L}\coloneqq\sum_{i=1}^{\operatorname{rank}\left({Y}\right)}u_{i}u_{i}^{*}$ and $P_{Y,R}\coloneqq\sum_{i=1}^{\operatorname{rank}\left({Y}\right)}v_{i}v_{i}^{*}$. Thus, $P_{Y,L}$ and $P_{Y,R}$ are the matrix representations in the standard bases of the projectors onto $\operatorname{ran}{Y}$ and $\ker{Y}^{\perp}$ respectively. It holds that $P_{Y,L}=YY^{\dagger}$ and $P_{Y,R}=Y^{\dagger}Y$. We note that the equality $P_{Y,L}=YY^{\dagger}$ is no longer true in general in a infinite-dimensional setting.
We let $\mathcal{H}$ and $\mathcal{K}$ be possibly infinite-dimensional separable Hilbert spaces over $\mathbb{C}$ or $\mathbb{R}$. By $\mathcal{B}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{00}(\mathcal{H}$, $\mathcal{K})$ and $L_{2}(\mathcal{H},\mathcal{K})$, we denote the linear spaces of bounded, compact, finite-rank and Hilbert–Schmidt operators between Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. The Hilbert–Schmidt norm of a linear operator between finite-dimensional spaces is equal to the Frobenius norm of any matrix representation of this operator in an orthonormal basis. By $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ we denote set of finite-rank operators, i.e. operators that are bounded, have finite-dimensional range, and have rank at most $r\in\mathbb{N}$. Further details of these classes of operators is given in Appendix A.
1.4 Outline
In Section 2, we recall the statement of [20, Theorem 2.1]. We then apply it to an elementary example in order to motivate the need for an alternative minimality property. In Section 3 we then formulate and prove a refinement of the statement of [20, Theorem 2.1] in the infinite-dimensional setting, with the modified minimality property. There, we also recover the modified finite-dimensional result of [20], characterise solutions that do not satisfy the minimality property, and construct approximate minimisers. In Section 4, it is shown that solutions may be unbounded. In Section 5, we discuss the cases in which solutions are bounded, and approximate unbounded solutions by bounded ones. We apply our results in a more concrete setting in Section 6, which includes examples from signal processing, linear operator learning and reduced rank regression. In Appendix A we recall certain concepts from functional analysis that we need in order to study the generalised rank-constrained approximation problem (1) in infinite dimensions. Finally, Appendix B contains some proofs of minor results used in the text.
2 Optimal approximation for matrices
In this section we recall the statement of [20, Theorem 2.1] and discuss the minimal norm property stated therein.
The statement of [20, Theorem 2.1], which provides an expression (3) for a solution of a rank-constrained approximation problem (4) and asserts that any such solution has a certain minimal norm property (5), is:
Claim of [20, Theorem 2.1]: Let $M\in\mathbb{C}^{m\times n}$, $B\in\mathbb{C}^{m\times p}$ and $C\in\mathbb{C}^{q\times n}$ be given. Then,
$\hat{X}=B^{\dagger}(P_{B,L}MP_{C,R})_{r}C^{\dagger},$ (3)
solves the low-rank approximation problem
$\min{\left\{\lVert M-BXC\rVert_{F},\ X\in\mathbb{C}^{q\times p},\ \operatorname{rank}\left({X}\right)\leq r\right\}},$ (4)
and satisfies the minimal norm property
[…]
A few remarks are in order. Firstly, in general not all solutions of (4) are of the form (3). Secondly, as $P_{C,R}=C^{\dagger}C$ and as in finite dimensions it holds that $P_{B,L}=BB^{\dagger}$, one may write (3) as $B^{\dagger}(BB^{\dagger}XC^{\dagger}C)_{r}C^{\dagger}$. Thirdly, also a necessary and sufficient condition for uniqueness of $\hat{X}$ is given in [20, Theorem 2.1]. Finally, if $(P_{B,L}MP_{C,R})_{r}$ is not uniquely specified, then the matrix $\hat{X}$ in (3) is not uniquely determined. Hence eq. 3 should be interpreted as a set of solutions. Furthermore, this lack of uniqueness extends also to $\lVert\hat{X}\rVert_{HS}$: if there exists more than one solution of the form (3), then these solutions might not all have the same norm. In other words, whether $\hat{X}$ in (3) is of minimal norm or not in general depends on the choice of truncated SVD $(P_{B,L}MP_{C,R})_{r}$. This fact indicates that the claim above is not sufficiently precise.
The following example demonstrates that if there is more than one solution to (4) of the form (3), then these solutions may have different Frobenius norm.
Example 2.1.
Let us take $n=m=2,\ p=q=3,\ r=1$, and
$M=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad B=C^{\top}=\begin{pmatrix}1&0&0\\ 0&\frac{1}{2}&0\\ \end{pmatrix}.$ (6)
Then,
$C^{\dagger}=(B^{\dagger})^{\top}=\begin{pmatrix}1&0&0\\ 0&2&0\\ \end{pmatrix},\quad P_{B,L}=P_{C,R}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
So $M=P_{B,L}P_{C,R}$ and its the rank-1 truncated SVDs are,
$M_{1,a}\coloneqq\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\quad M_{1,b}\coloneqq\begin{pmatrix}0&0\\ 0&1\\ \end{pmatrix}.$
The two minimizing matrices of the form (3) then are,
$\hat{X}_{a}\coloneqq\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\\ \end{pmatrix},\quad\hat{X}_{b}\coloneqq\begin{pmatrix}0&0&0\\ 0&4&0\\ 0&0&0\\ \end{pmatrix},$
However, the Frobenius norms of $\hat{X}_{a},\hat{X}_{b}$ differ, and thus cannot both be minimal. In particular, (5) does not hold.
The preceding example suggests that [20, Theorem 2.1] can be made more precise, either by removing the minimum norm condition (5), or by replacing this condition by a weaker condition which all solutions $\hat{X}$ given by (3) satisfy. Such a weaker condition must thus be independent of the choice of truncated SVD $(P_{B,L}MP_{C,R})_{r}$. Furthermore, if $C$ is a linear operator mapping between infinite-dimensional spaces, then ${C}^{\dagger}$ will in general not be bounded (cf. Proposition A.5), and thus cannot be Hilbert–Schmidt. It can then happen that also solutions $\hat{X}$ of the form (3) are not Hilbert–Schmidt, in which case $\lVert\hat{X}\rVert_{HS}$ may be ill-defined, and we need a minimality condition that does not involve the Frobenius norm in order to generalise this condition to the infinite-dimensional setting. For these two reasons we motivate an alternative condition, by continuing our Example 2.1.
Example 2.1 (Continued).
Let us put
[…]
with $X_{a,11}=\big{(}\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\big{)}$ and $X_{b,11}=\big{(}\begin{smallmatrix}0&0\\ 0&4\end{smallmatrix}\big{)}$, so $\hat{X}_{a}(0)=\hat{X}_{a}$ and $\hat{X}_{b}(0)=\hat{X}_{b}$. For $c\in\{a,b\}$, we notice that $\hat{X}_{c}(\alpha)$ also solves the minimisation problem (4), for any $\alpha=(\alpha_{i})_{i=1}^{5}\in\mathbb{C}^{5}$, as $B\hat{X}_{a}(\alpha)C=\left(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\right)$ and $B\hat{X}_{b}(\alpha)C=\left(\begin{smallmatrix}0&0\\ 0&1\end{smallmatrix}\right)$. Hence, $B\hat{X}_{c}(\alpha)C=B\hat{X}_{c}(0)C$ for all $\alpha$. Furthermore, any solution to (4) can be written as $\hat{X}_{c}(\alpha)$ for some $\alpha\in\mathbb{C}^{5}$ and $c\in\{a,b\}$. To see this, let us write an element of $\mathbb{C}^{3\times 3}$ as, for $x_{11},x_{12},x_{21},x_{22}\in\mathbb{C}$,
$X(\alpha)\coloneqq\begin{pmatrix}x_{11}&x_{12}&\alpha_{1}\\ x_{21}&x_{22}&\alpha_{2}\\ \alpha_{3}&\alpha_{4}&\alpha_{5}\end{pmatrix}.$
Like $B\hat{X}_{c}(\alpha)C$, also $BX(\alpha)C$ does not depend on $\alpha$. By the Eckart–Young Theorem (c.f. Proposition A.7), for $X(\alpha)$ to be a solution of (4), we need $BX(\alpha)C$ to be a rank-1 SVD of $M$. That is, we need $x_{12}=x_{21}=0$ and either $x_{11}=1$ or $x_{22}=4$.
Thus, among all the solutions of (4), i.e. among all $\hat{X}_{c}(\alpha)$, only $\hat{X}_{a}(0)$ and $\hat{X}_{b}(0)$ are of the form (3), and only $\hat{X}_{a}(0)$ satisfies (5). For this example, we now construct an alternative to (5), which both $\hat{X}_{a}(0)$ and $\hat{X}_{b}(0)$ satisfy and which can still be interpreted as a minimality condition.
Let $e_{i}$ denote the $i$-th canonical basis vector of $\mathbb{C}^{3}$ and $V\coloneqq\operatorname{span}{\left(e_{1},e_{2}\right)}\subset\mathbb{C}^{3}$. Let $P_{V}$ the matrix representation of the projector onto $V$ in the canonical basis of $\mathbb{C}^{3}$. Then for $B,C$ defined in (6), $P_{B,R}=P_{V}=P_{C,L}$. We observe $P_{B,L}\hat{X}(\alpha)P_{C,R}=P_{V}\hat{X}_{c}(\alpha)P_{V}=\hat{X}_{c}(0)$ for any $\alpha$, and compute,
$\lVert\hat{X}_{c}(\alpha)\rVert_{HS}^{2}=\lVert X_{c,11}\rVert_{HS}^{2}+\sum_{i=1}^{5}\lvert\alpha_{i}\rvert^{2}\geq\lVert X_{c,11}\rVert_{HS}^{2}=\lVert\hat{X}_{c}(0)\rVert_{HS}^{2}.$ (7)
Hence, for fixed $c\in\{a,b\}$, among the solutions $\hat{X}_{c}(\alpha)$, only $\hat{X}_{c}(0)$ has minimal Hilbert–Schmidt norm. We recall that the $\hat{X}_{c}(\alpha)$ satisfy $P_{V}\hat{X}_{c}(\alpha)P_{V}=\hat{X}_{c}(\alpha)$ if and only if $\alpha=0$. That is, it is precisely the solutions $\hat{X}$ of (4) which satisfy
$\hat{X}=P_{B,R}\hat{X}P_{C,L},$ (8)
that are the minimisers in (7) for some $c\in\{a,b\}$. Hence (8) provides an alternative for the minimal norm property (5). So, while $X_{a}(0)$ and $X_{b}(0)$ might not have the same norm, $X_{c}(0)$ has minimal norm among all $X_{c}(\alpha),\alpha\in\mathbb{C}^{5}$, for fixed $c\in\{a,b\}$, and $X_{a}(0),X_{b}(0)$ are in turn those solutions of (4) which satisfy (8).
To generalise the reasoning of Example 2.1 as to why (8) can be interpreted as a minimality condition, notice that (8) is equivalent to $(I-P_{B,R})\hat{X}P_{C,L}=0$ and $\hat{X}(I-P_{C,L})=0$, since any $X\in\mathbb{C}^{q\times p}$ can be decomposed as
$X=P_{B,R}XP_{C,L}+(I-P_{B,R})XP_{C,L}+X(I-P_{C,L}).$
Now, $(I-P_{B,R})\hat{X}P_{C,L}=0$ is equivalent to the property that $\hat{X}$ maps $\operatorname{ran}{C}$ into $\ker{B}^{\perp}$ and $\hat{X}(I-P_{C,L})=0$ is equivalent to the property that $\hat{X}$ maps $\operatorname{ran}{C}^{\perp}$ to $\{0\}$. In other words, if we express $\hat{X}\in\mathbb{C}^{p\times q}$ relative to the decompositions $\operatorname{ran}{C}\oplus\operatorname{ran}{C}^{\perp}$ of its domain and $\ker{B}^{\perp}\oplus\ker{B}$ of its range,
$\hat{X}=\begin{pmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{pmatrix},$
with $X_{11}\in\mathbb{C}^{s\times t}$, $X_{12}\in\mathbb{C}^{s\times(q-t)}$, $X_{21}\in\mathbb{C}^{(p-s)\times t}$, $X_{22}\in\mathbb{C}^{(p-s)\times(q-t)}$, $t\coloneqq\dim\operatorname{ran}{C}$, $s\coloneqq\dim\ker{B}^{\perp}$, then $X_{21}$, $X_{12}$, $X_{22}$ are all zero. If (8) fails to hold, then at least one of the blocks $X_{12},X_{21},X_{22}$ must be nonzero, and the Hilbert–Schmidt norm of $\hat{X}$ is larger than that of $X_{11}$. Thus, we can view (8) as expressing a minimality property, without referring to a norm.
If $\hat{X}$ in (3) is unique, then $\lVert\hat{X}\rVert_{HS}$ is indeed minimal among all minimisers $X$ of (4), in which case the statement of [20, Theorem 2.1] is still valid. Of course, this holds only when Hilbert–Schmidt norms can in fact be computed, which is no longer true in general for infinite-dimensional spaces. In the next section we introduce a statement similar to [20, Theorem 2.1] with the modified minimality property (8) in an infinite-dimensional setting.
3 Optimal approximations for Hilbert–Schmidt operators on separable Hilbert spaces
We are now ready to formulate (3) and (4) with the minimality property (8) in infinite dimensions. In Section 3.1, we give necessary and sufficient conditions for a solution to exist and describe the solution. An interpretation of the minimality property is also given, as well as the finite-dimensional modification of [20, Theorem 2.1]. When no solution exists, i.e. when the minimum Hilbert–Schmidt approximation error is not attained, one can still explicitly construct arbitrarily good approximations, as is shown in Section 3.2. When a solution does exist, the corresponding optimal Hilbert–Schmidt approximation error is computed in Section 3.3.
3.1 Problem formulation and solution
As in the finite-dimensional setting of Section 2, we again need the orthogonal projectors $P_{\overline{\operatorname{ran}{A}}}$ and $P_{\ker{A}^{\perp}}$ onto the closure of the range and orthogonal complement of the kernel of a bounded linear operator $A$. We note by (36) that $P_{\ker{A}^{\perp}}=A^{\dagger}A$ as in the finite-dimensional setting. However, unlike the finite-dimensional setting, it holds by (37) that $P_{\overline{\operatorname{ran}{A}}}$ merely extends $AA^{\dagger}$, as $AA^{\dagger}$ is only defined on the dense subspace $\operatorname{dom}{A^{\dagger}}$, whereas $P_{\overline{\operatorname{ran}{A}}}$ is defined everywhere.
Setting 3.1.
Let $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3},\mathcal{H}_{4}$ be separable Hilbert spaces, let $r\in\mathbb{N}$ and let $M\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$, $B\in\mathcal{B}(\mathcal{H}_{3},\mathcal{H}_{4})$ and $C\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$ be given. We define
$\mathcal{X}_{r}$$\coloneqq\left\{X:\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},\ \operatorname{dom}{X}\supset\operatorname{dom}{C^{\dagger}},\right.$
$\left.\qquad\dim\operatorname{ran}{X}\leq r,BXC\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})\right\},$
$\mathcal{Y}_{r}$$\coloneqq\left\{Y\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4}):\ \forall\tilde{Y}\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\right.$
$\left.\qquad\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}\leq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-\tilde{Y}\rVert_{HS},\right.$
$\left.\qquad\operatorname{ran}{Y}\subset\operatorname{dom}{B^{\dagger}}\right\}.$
Thus in 3.1, $\mathcal{X}_{r}$ contains in particular all operators of rank at most $r$, which are bounded by definition. We stress that $\mathcal{X}_{r}$ in general also contains unbounded operators whose range has dimension at most $r$. Furthermore, $\mathcal{Y}_{r}$ is the set of solutions of the reduced rank problem
$\min\{\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}:\ Y\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\},$ (9)
that also map into $\operatorname{dom}{B^{\dagger}}$. Hence Proposition A.7, $\mathcal{Y}_{r}$ consists of all rank-$r$ truncated SVDs of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$, as defined in (38), that map into $\operatorname{dom}{B^{\dagger}}$.
We can now state the infinite-dimensional problem and its solution.
Theorem 3.2.
Assume 3.1. There exists a solution to the problem
$\min\{\lVert M-BXC\rVert_{HS},\ X\in\mathcal{X}_{r}\},$ (10)
if and only if $\mathcal{Y}_{r}\not=\emptyset$. In that case, for each $(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}\in\mathcal{Y}_{r}$ a solution is given by
$\hat{X}:\operatorname{dom}{C^{\dagger}}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},\quad\hat{X}=B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger},$ (11)
Furthermore, $\hat{X}$ satisfies
$X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\text{on }\operatorname{dom}{X}.$ (12)
This solution is unique, in the sense that $X=\tilde{X}$ on $\operatorname{dom}{C^{\dagger}}$ for solutions $X,\tilde{X}$ of the form (11) satisfying (12), if and only if one of the following conditions holds:
  1. (a)
    $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4}),$
  2. (b)
    $\sigma_{r}({P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}})>\sigma_{r+1}({P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}}).$
The conditions (a) or (b) ensure the rank-$r$ reduced SVD in (11) is unique, c.f. Section A.3. Furthermore, if $\operatorname{ran}{M}\subset\operatorname{dom}{B}^{\dagger}$, then $P_{\overline{\operatorname{ran}{B}}}M=BB^{\dagger}M$ by (37) and
$\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}}\subset\operatorname{ran}{P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}}\subset\operatorname{ran}{B}.$
The condition $\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}}\subset\operatorname{dom}{B^{\dagger}}$ is then satisfied.
Remark 3.3 (Solution domain).
In general, not all solutions of (10) that satisfy (12) are given by (11), as they might have a domain larger than $\operatorname{dom}{C^{\dagger}}=\operatorname{dom}{X^{\dagger}}$, but they agree with $\hat{X}$ on $\operatorname{dom}{C^{\dagger}}$, i.e. they extend $\hat{X}$.
Remark 3.4 (Equivalent domain condition).
Since $\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}}\subset\operatorname{ran}{P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}}\subset\overline{\operatorname{ran}{B}}$ and $\operatorname{dom}{B}^{\dagger}=\operatorname{ran}{B}\oplus\operatorname{ran}{B}^{\perp}$, the requirement $\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}}\subset\operatorname{dom}{B}^{\dagger}$ in the definition of $\mathcal{Y}_{r}$ is equivalent to $\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}}\subset\operatorname{ran}{B}$.
Remark 3.5 (Case with $B$, $C$ having closed range).
By Proposition A.5 item (iv),
  1. (i)
    If $C$ has closed range, then $\mathcal{X}_{r}=B_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ and $P_{\overline{\operatorname{ran}{C}}}=CC^{\dagger}$. In this case, all solutions of (10) that satisfy (12) are given by (11).
  2. (ii)
    If $B$ has closed range, then $\operatorname{dom}{B^{\dagger}}=\mathcal{H}_{4}$, so $\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}}\subset\operatorname{dom}{B^{\dagger}}$ trivially, hence $\mathcal{Y}_{r}=\{(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}\}$ and so $\mathcal{Y}_{r}$ is nonempty.
In order to prove the theorem, we define the following solution spaces:
$\mathcal{S}_{r}$$\coloneqq\left\{X\in\mathcal{X}_{r}:\ \lVert M-BXC\rVert_{HS}\leq\lVert M-B\tilde{X}C\rVert_{HS}\ \forall\tilde{X}\in\mathcal{X}_{r}\right\},$
$\mathcal{S}_{r,*}$$\coloneqq\left\{X\in\mathcal{S}_{r}:\ X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\text{ on }\operatorname{dom}{X}\right\},$
Thus, $\mathcal{S}_{r}$ is the set of solutions to the reduced rank problem (10). Further, $\mathcal{S}_{r,*}$ is the set of solutions to (10) that also satisfy (12). We must show, under the existence condition $\mathcal{Y}_{r}\not=\emptyset$, that $\hat{X}$ defined in (11) lies in $\mathcal{S}_{r,*}$ and, under the uniqueness condition (a) or (b), that $X=\tilde{X}$ on $\operatorname{dom}{C^{\dagger}}$ for any $X,\tilde{X}\in\mathcal{S}_{r,*}$. To this end, we show that solutions to (10) and (12) can be constructed by a solution to (9) which maps into $\operatorname{dom}{B^{\dagger}}$ and vice versa, thereby drawing a connection between the sets $\mathcal{S}_{r,*}$ and $\mathcal{Y}_{r}$. The following lemma is crucial for this purpose.
Lemma 3.6.
Let $M$, $B$, $C$ and $\mathcal{X}_{r}$ be as in Theorem 3.2 and $c\coloneqq\lVert M\rVert_{HS}^{2}-\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\rVert_{HS}^{2}$. For any $X\in\mathcal{X}_{r}$,
$\lVert M-BXC\rVert_{HS}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+c.$ (13)
Proof of Lemma 3.6.
For any separable Hilbert space $\mathcal{K}$ with ONB $(e_{i})_{i}$, we have,
$\langle P_{\overline{\operatorname{ran}{B}}}S,(I-P_{\overline{\operatorname{ran}{B}}})T\rangle_{HS}=\sum_{i}\langle P_{\overline{\operatorname{ran}{B}}}Se_{i},(I-P_{\overline{\operatorname{ran}{B}}})Te_{i}\rangle=0,\quad T,S\in\mathcal{B}(\mathcal{K},\mathcal{H}_{4}),$
$\langle SP_{\ker{C}^{\perp}},T(I-P_{\ker{C}^{\perp}})\rangle_{HS}=\langle(I-P_{\ker{C}^{\perp}})T^{*},P_{\ker{C}^{\perp}}S^{*}\rangle_{HS}=0,\quad T,S\in\mathcal{B}(\mathcal{H}_{1},\mathcal{K}).$
Furthermore, the Moore–Penrose equations (34), (36) and (37) in Proposition A.5 state that $P_{\overline{\operatorname{ran}{B}}}B=BB^{\dagger}B=B$ and $CP_{\ker{C}^{\perp}}=CC^{\dagger}C=C$. Therefore, for any $X\in\mathcal{X}_{r}$, we have $P_{\overline{\operatorname{ran}{B}}}BXCP_{\ker{C}^{\perp}}=BXC\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ and,
$\lVert M-BXC\rVert_{HS}^{2}$ $=$ $\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+\lVert(I-P_{\overline{\operatorname{ran}{B}}})MP_{\ker{C}^{\perp}}\rVert_{HS}^{2}$
$+\lVert P_{\overline{\operatorname{ran}{B}}}M(I-P_{\ker{C}^{\perp}})\rVert_{HS}^{2}+\lVert(I-P_{\overline{\operatorname{ran}{B}}})M(I-P_{\ker{C}^{\perp}})\rVert_{HS}^{2}$
$=$ $\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+c.$
Using Lemma 3.6, we can now relate the sets $\mathcal{S}_{r}$ and $\mathcal{Y}_{r}$.
Lemma 3.7.
Let $M,B,C$ and $\mathcal{X}_{r}$ be as in Theorem 3.2. Then
  1. (i)
    For $Y\in\mathcal{Y}_{r}$ it holds that ${B^{\dagger}YC^{\dagger}}$ is a well-defined linear operator on $\operatorname{dom}{C^{\dagger}}$, which satisfies $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$ and $BB^{\dagger}YC^{\dagger}C=Y$,
  2. (ii)
    For $X\in\mathcal{X}_{r}$, $BXC\in\mathcal{Y}_{r}$ if and only if $X\in\mathcal{S}_{r}$,
  3. (iii)
    For $Y\in\mathcal{Y}_{r}$ it holds that $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r,*}$,
  4. (iv)
    For $X\in\mathcal{S}_{r,*}$ there exists $Y\in\mathcal{Y}_{r}$ such that $X=B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$.
If we define the map $j:\mathcal{X}_{r}\rightarrow\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})$, $X\mapsto BXC$, then the “if” part of (ii) in Lemma 3.7 implies that $j(\mathcal{S}_{r})\subset\mathcal{Y}_{r}$ and the “only if” part that $j(\mathcal{X}_{r}\backslash\mathcal{S}_{r})\subset\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})\backslash\mathcal{Y}_{r}$. Item (i) states that ${\left.\kern-1.2ptj\vphantom{\big{|}}\right|_{\mathcal{S}_{r}}}$ maps onto $\mathcal{Y}_{r}$, hence has a right inverse, and that a right inverse is given by the map $Y\mapsto B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$. By item (iii), the image of this right inverse maps into $\mathcal{S}_{r,*}$. By item (iv), every $X\in\mathcal{S}_{r,*}$ agrees with $B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$, for some $Y\in\mathcal{Y}_{r}$.
Proof of Lemma 3.7.
(i) Let $Y\in\mathcal{Y}_{r}$. That is, let $Y$ be a rank-$r$ truncated SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ satisfying $\operatorname{ran}{Y}\subset\operatorname{dom}{B^{\dagger}}$. It follows that $B^{\dagger}Y$ is well-defined on all of $\mathcal{H}_{2}$. By Lemma A.1, $B^{\dagger}Y$ is in fact bounded and of rank at most $r$. Thus, by the definition of compositions of linear maps in Section A.1,
$\operatorname{dom}{B^{\dagger}YC^{\dagger}}=(C^{\dagger})^{-1}\left(\operatorname{dom}{B^{\dagger}Y}\right)=\operatorname{dom}{C^{\dagger}},$
and
$\dim\operatorname{ran}{B^{\dagger}YC^{\dagger}}\leq\operatorname{rank}\left({Y}\right)\leq r.$
To show that $B^{\dagger}YC^{\dagger}\in\mathcal{X}_{r}$, it remains to show $B(B^{\dagger}YC^{\dagger})C\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Now, $BB^{\dagger}$ is the identity on $\operatorname{ran}{B}$ by (37). As $\operatorname{ran}{Y}\subset\operatorname{ran}{B}$, it holds that $BB^{\dagger}Y=Y$. Further,
$\ker{Y}\supset\ker{P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}}\supset\ker{P_{\ker{C}^{\perp}}}=\ker{C},$
where the first inclusion holds because $Y$ is a truncated SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. The inclusion $\ker{Y}\supset\ker{C}$ and (36) imply $Y=YP_{\ker{Y}^{\perp}}=YP_{\ker{C}^{\perp}}=YC^{\dagger}C$. We conclude,
$B(B^{\dagger}YC^{\dagger})C=YC^{\dagger}C=Y.$ (14)
In particular, $B(B^{\dagger}YC^{\dagger})C$ has rank at most $r$ and is hence Hilbert–Schmidt. Thus, $B^{\dagger}YC^{\dagger}\in\mathcal{X}_{r}$. Finally, $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$, since for any $\tilde{X}\in\mathcal{X}_{r}$ it holds by applying (13) twice, (14) and the definition of $Y$,
$\lVert M-B\tilde{X}C\rVert_{HS}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\tilde{X}C\rVert_{HS}^{2}+c\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}^{2}+c$
$=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BB^{\dagger}YC^{\dagger}C\rVert_{HS}^{2}+c=\lVert M-B(B^{\dagger}YC^{\dagger})C\rVert_{HS}^{2}.$
(ii) Let $X\in\mathcal{X}_{r}$, and let us assume that $BXC\in\mathcal{Y}_{r}$. By applying (13) twice and by the definition of $\mathcal{Y}_{r}$, we have for all $\tilde{X}\in\mathcal{X}_{r}$,
$\lVert M-B\tilde{X}C\rVert_{HS}^{2}$ $=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\tilde{X}C\rVert_{HS}^{2}+c$
$\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B{X}C\rVert_{HS}^{2}+c=\lVert M-B{X}C\rVert_{HS}^{2}.$
Hence, $X\in\mathcal{S}_{r}$. Conversely, let us assume $X\in\mathcal{S}_{r}$. Let $Y\in\mathcal{Y}_{r}$. By item (i), $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$. Using (13) twice and item (i), we find for all $\tilde{Y}\in\mathcal{B}_{00,r}(\mathcal{H}_{1},\mathcal{H}_{4})$,
$\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-\tilde{Y}\rVert_{HS}^{2}$ $\geq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}^{2}=\lVert M-B(B^{\dagger}YC^{\dagger})C\rVert_{HS}^{2}-c$
$\geq\lVert M-BXC\rVert_{HS}^{2}-c=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS},$
where we used $Y\in\mathcal{Y}_{r}$ for the first inequality and $X\in\mathcal{S}_{r}$ for the second inequality. Thus $BXC\in\mathcal{Y}_{r}$.
(iii) Let $Y\in\mathcal{Y}_{r}$. By item (i), $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r}$. By (35),
$P_{\ker{B}^{\perp}}B^{\dagger}YC^{\dagger}P_{\overline{\operatorname{ran}{C}}}=B^{\dagger}BB^{\dagger}YC^{\dagger}CC^{\dagger}=B^{\dagger}YC^{\dagger}\quad\text{on }\operatorname{dom}{C^{\dagger}}.$
Thus, $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r,*}$.
(iv) Let $X\in\mathcal{S}_{r,*}$, so $Y\coloneqq BXC\in\mathcal{Y}_{r}$ by item (ii). By definition of $\mathcal{S}_{r,*}$ it holds that $X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}$ on $\operatorname{dom}{C^{\dagger}}\subset\operatorname{dom}{X}$. Hence, using (36) and (37), $X=B^{\dagger}(BXC)C^{\dagger}=B^{\dagger}YC^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$.
Proof of Theorem 3.2.
Suppose that $\mathcal{Y}_{r}\not=\emptyset$ and let $Y\in\mathcal{Y}_{r}$. Then Lemma 3.7 item (iii) gives that $B^{\dagger}YC^{\dagger}\in\mathcal{S}_{r,*}$, i.e. (11) solves (10) and satisifes (12). Conversely, suppose that (10) admits a solution $\hat{X}\in\mathcal{X}_{r}$, so $\hat{X}\in\mathcal{S}_{r}$. By Lemma 3.7 item (ii) and Proposition A.7, $Y\coloneqq B\hat{X}C$ is a truncated rank-$r$ SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. But $\operatorname{ran}{Y}\subset\operatorname{ran}{B}\subset\operatorname{dom}{B^{\dagger}}$, so $Y\in\mathcal{Y}_{r}$.
It remains to show that all elements of $S_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$ under conditions (a) or (b). By the construction of the SVD in Section A.3, $\mathcal{Y}_{r}$ contains at most one element if and only if (a) or (b) holds. In this case, by item (iv) of Lemma 3.7, all elements of $\mathcal{S}_{r,*}$ agree on $\operatorname{dom}{C^{\dagger}}$. This concludes the proof of uniqueness. ∎
As a corollary to Theorem 3.2 we obtain the result of [20, Theorem 2.1] but with the modified minimality property. We recall that $P_{A,L}$ and $P_{A,R}$ are defined in Section 2 and that $P_{A,L}=AA^{\dagger}$ and $P_{A,R}=A^{\dagger}A$.
Corollary 3.8.
Let $B\in\mathbb{C}^{m\times p},C\in\mathbb{C}^{q\times n},M\in\mathbb{C}^{m\times n}.$ Then
$\hat{X}=B^{\dagger}(P_{B,L}MP_{C,R})_{r}C^{\dagger}$
is a solution to the problem
$\min\{\lVert M-BXC\rVert_{F},\ X\in\mathbb{C}^{p\times q},\operatorname{rank}\left({X}\right)\leq r\}.$
Furthermore, it satisfies
$X=P_{B,R}XP_{C,L}.$ (15)
This solution is unique if and only if one of the following conditions holds,
  1. (a)
    $r\geq\operatorname{rank}\left({P_{B,L}MP_{C,R}}\right),$
  2. (b)
    $\sigma_{r}({P_{B,L}MP_{C,R}})>\sigma_{r+1}({P_{B,L}MP_{C,R}}).$
Proof.
We associate a matrix $A\in\mathbb{C}^{t\times s}$ with the linear operator $L_{A}\in\mathcal{B}(\mathbb{C}^{s},\mathbb{C}^{t})$ that left multiplies $x\in\mathbb{C}^{s}$ with $A$, i.e $L_{A}x=Ax$. Hence $A$ is the expression of $L_{A}$ in the orthogonal standard bases of $\mathbb{C}^{s}$ and $\mathbb{C}^{t}$, so we have the equality of norms $\lVert L_{A}\rVert_{HS}=\lVert A\rVert_{F}$. Furthermore, one may check that $L_{A}^{\dagger}=L_{A^{\dagger}}$, $(L_{A})_{r}=L_{(A)_{r}}$, $P_{\overline{\operatorname{ran}{L_{A}}}}=P_{\operatorname{ran}{L_{A}}}=P_{\operatorname{ran}{A}}=L_{P_{A,L}}$ and $P_{\ker{L_{A}}^{\perp}}=P_{\ker{A}^{\perp}}=L_{P_{A,R}}$. As all finite-dimensional linear maps are Hilbert–Schmidt operators and have closed range, Remark 3.5 implies that we can apply Theorem 3.2 to $L_{M},L_{B},L_{C}$ with $\mathcal{X}_{r}=\mathcal{B}_{00,r}(\mathbb{C}^{q},\mathbb{C}^{p})\simeq\{X\in\mathbb{C}^{p\times q},\operatorname{rank}\left({X}\right)\leq r\}$ and where $\mathcal{Y}_{r}$ consists of the truncated rank-$r$ SVDs of $P_{{\operatorname{ran}{B}}}L_{M}P_{\ker{C}^{\perp}}$. The result now follows from the conclusions of Theorem 3.2 by expressing the linear operators again as matrices. ∎
Theorem 3.2 generalises the statement in [20] about existence and uniqueness of $\hat{X}$ of the optimal low-rank approximation problem (4)–(5) in the finite-dimensional setting to the setting of possibly infinite-dimensional separable Hilbert spaces. The minimality property (5) in [20, Theorem 2.1] has been replaced by (12) in Theorem 3.2.
We now present an interpretation of (8) that parallels the interpretation of property (5) in the finite-dimensional setting given at the end of Section 2. Let us take $X\in\mathcal{X}_{r}$. Then $\operatorname{dom}{X}\supset\operatorname{ran}{C}$, so $\mathcal{D}\coloneqq\operatorname{dom}{X}\cap\overline{\operatorname{ran}{C}}$ contains $\operatorname{ran}{C}$. Therefore, $\mathcal{D}\subset\overline{\operatorname{ran}{C}}$ densely and $\mathcal{D}^{\perp}=\operatorname{ran}{C}^{\perp}$. Also, $\operatorname{dom}{X}=\mathcal{D}\oplus\operatorname{ran}{C}^{\perp}$. Indeed, when $h\in\operatorname{dom}{X}\subset\mathcal{H}_{2}=\overline{\operatorname{ran}{C}}\oplus\operatorname{ran}{C}^{\perp}$, then $h=h_{1}+h_{2}$ for some $h_{1}\in\overline{\operatorname{ran}{C}},h_{2}\in\operatorname{ran}{C}^{\perp}\subset\operatorname{dom}{C^{\dagger}}\subset\operatorname{dom}{X}$, hence $h_{1}=h-h_{2}\in\operatorname{dom}{X}$.
Let us decompose $\mathcal{H}_{3}=\ker{B}^{\perp}\oplus\ker{B}$ and $\operatorname{dom}{X}=\mathcal{D}\oplus\operatorname{ran}{C}^{\perp}$ We define,
$X_{11}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}\right)\vphantom{\big{|}}\right|_{\mathcal{D}}}:\mathcal{D}\subset\overline{\operatorname{ran}{C}}\rightarrow\ker{B}^{\perp}$
$X_{12}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}^{\perp}}XP_{\operatorname{ran}{C}^{\perp}}\right)\vphantom{\big{|}}\right|_{\operatorname{ran}{C}^{\perp}}}:\operatorname{ran}{C}^{\perp}\rightarrow\ker{B}^{\perp}$
$X_{21}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}\right)\vphantom{\big{|}}\right|_{\mathcal{D}}}:\mathcal{D}\subset\overline{\operatorname{ran}{C}}\rightarrow\ker{B}$
$X_{22}$ $\coloneqq{\left.\kern-1.2pt\left(P_{\ker{B}}XP_{\operatorname{ran}{C}^{\perp}}\right)\vphantom{\big{|}}\right|_{\operatorname{ran}{C}^{\perp}}}:\operatorname{ran}{C}^{\perp}\rightarrow\ker{B}.$
While $X_{12},X_{22}$ are defined everywhere on $\operatorname{ran}{C}^{\perp}$, $X_{11},X_{21}$ are densely defined in $\overline{\operatorname{ran}{C}}$, as $\mathcal{D}\subset\overline{\operatorname{ran}{C}}$ densely. By Proposition A.5 item (iv), $X_{11},X_{21}$ are everywhere defined when $\operatorname{ran}{C}$ is closed. We can now write $X$ relative to these decompositions of $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ as
$X=\begin{pmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{pmatrix},$ (16)
with the understanding that for $h_{1}+h_{2}\in\mathcal{D}\oplus\operatorname{ran}{C}^{\perp}=\operatorname{dom}{X}$,
$X(h_{1},h_{2})$ $=\begin{pmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{pmatrix}\begin{pmatrix}h_{1}\\ h_{2}\end{pmatrix}$
$=(X_{11}h_{1}+X_{12}h_{2},X_{21}h_{1}+X_{22}h_{2})\in\ker{B}^{\perp}\oplus\ker{B}.$
Since
$X=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}+(I-P_{\ker{B}^{\perp}})XP_{\overline{\operatorname{ran}{C}}}+X(I-P_{\overline{\operatorname{ran}{C}}}),$ (17)
the minimality condition (12) is equivalent to both $X(I-P_{\overline{\operatorname{ran}{C}}})=0$ and $(I-P_{\ker{B}^{\perp}})XP_{\overline{\operatorname{ran}{C}}}=0$ on $\operatorname{dom}{X}$. It follows, respectively, that $X_{12}=X_{22}=0$ and that $X_{21}=0$. This generalises the interpretation of (8) in Section 2 to infinite-dimensional separable Hilbert spaces: the minimality of $\hat{X}$ of the form (11) is in the sense that all blocks in (16) except for the upper left block $X_{11}$ are zero.
As noted in Remark 3.3, we have only characterised the solutions of (10) that satisfy (12) on $\operatorname{dom}{C}^{\dagger}$. That is, we have characterised $\mathcal{S}_{r,*}\cap\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\}$. We can now charachterise all solutions of (10), that need not satisfy the minimality property (12), on $\operatorname{dom}{C}^{\dagger}$.
Corollary 3.9.
The problem (10) admits a solution if and only if $\mathcal{Y}_{r}\not=\emptyset$. In that case,
$\mathcal{S}_{r}\cap\left\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\right\}=\left\{\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}}:\ \hat{X}\in\mathcal{S}_{r,*},\right.$
$\left.T:\operatorname{dom}{C}^{\dagger}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},S:\operatorname{ran}{C}^{\perp}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}\right\}.$
Proof.
It follows directly from Theorem 3.2 that (10) has a solution if and only if $\mathcal{Y}_{r}\not=\emptyset$. Now let $X=\hat{X}+SP_{\operatorname{ran}{C}^{\perp}}+P_{\ker{B}}T$ with $\hat{X}$, $S$ and $T$ as specified. Then $\operatorname{dom}{X}=\operatorname{dom}{\hat{X}}=\operatorname{dom}{C}^{\dagger}$ and $BXC=B\hat{X}C$. But $\hat{X}\in\mathcal{S}_{r}$, so $X\in\mathcal{S}_{r}$. This shows the $\supset$” inclusion holds. Conversely, let $X\in\mathcal{S}_{r}$ with $\operatorname{dom}{X}=\operatorname{dom}{C^{\dagger}}$. We define $\hat{X}=B^{\dagger}BXCC^{\dagger}$, so $\hat{X}\in\mathcal{S}_{r,*}$ by Lemma 3.7 Items (ii) and (iii). In particular, using also (36) and (37), we have on $\operatorname{dom}{C}^{\dagger}=\operatorname{dom}{X}$ that $P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}=B^{\dagger}BXCC^{\dagger}=\hat{X}$. We also define $S\coloneqq XP_{\operatorname{ran}{C}^{\perp}}$ and $T\coloneqq{\left.\kern-1.2ptP_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{C}^{\dagger}}}$. So $S$ and $T$ well-defined, as $P_{\operatorname{ran}{C}^{\perp}}$ and ${\left.\kern-1.2ptP_{\overline{\operatorname{ran}{C}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{C}^{\dagger}}}$ both map into $\operatorname{dom}{C}^{\dagger}=\operatorname{dom}{X}$. Furthermore, $S=SP_{\operatorname{ran}{C}^{\perp}}$ and $T=P_{\ker{B}}T$. Hence using (17):
$X$ $=P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}+P_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}+XP_{\operatorname{ran}{C}^{\perp}}$
$=\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}},$
where the second equality only holds on $\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}$. ∎
We end this section by noting that the solutions of (10) are all unbounded when only the solutions that also satisfy the minimality property (12) are unbounded.
Corollary 3.10.
If $\mathcal{S}_{r,*}$ contains only unbounded solutions, then all solutions of problem (10) are unbounded.
Proof.
Let $\tilde{X}$ be a solution of (10). By Corollary 3.9, we have that $\mathcal{S}_{r,*}\not=\emptyset$ and on $\operatorname{dom}{C^{\dagger}}$ that $\tilde{X}=\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}}$ for some $\hat{X}\in\mathcal{S}_{r,*}$, $T:\operatorname{dom}{C}^{\dagger}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}$ and $S:\operatorname{ran}{C}^{\perp}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}$. We have $\hat{X}=P_{\ker{B}^{\perp}}\hat{X}$ by (12). For $h\in\operatorname{ran}{C}$ it holds that $P_{\operatorname{ran}{C}^{\perp}}h=0$, hence,
$\lVert\tilde{X}h\rVert^{2}=\lVert\hat{X}h+P_{\ker{B}}Th\rVert^{2}=\lVert\hat{X}h\rVert^{2}+\lVert P_{\ker{B}}Th\rVert^{2}\geq\lVert\hat{X}h\rVert^{2}.$
Let us assume $\hat{X}$ is unbounded. Then there exists a bounded sequence $(h_{n})_{n}$ in $\operatorname{dom}{C^{\dagger}}$ satisfying $\lVert\hat{X}h_{n}\rVert\rightarrow\infty$. Since $\hat{X}$ satisfies (12), $\hat{X}h_{n}=\hat{X}P_{\operatorname{ran}{C}}h_{n}$. So $k_{n}\coloneqq P_{\operatorname{ran}{C}}h_{n}\in\operatorname{ran}{C}$ is a bounded sequence and $\lVert\hat{X}k_{n}\rVert$ diverges. But then also $\lVert\tilde{X}k_{n}\rVert$ diverges, showing that $\tilde{X}$ is unbounded. ∎
3.2 The case of no solutions and approximate minimisers
Even when the condition $\mathcal{Y}_{r}\not=\emptyset$ with $\mathcal{Y}_{r}$ defined in 3.1 fails to hold, that is, even when $\operatorname{ran}{Y}\subset\operatorname{dom}{B}^{\dagger}$ does not hold for any rank-$r$ reduced SVD $Y$ of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$, one can still construct a minimizing sequence for (10) that satisfies (12). To do so, suppose $Y\coloneqq(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ has SVD $\sum_{i=1}^{r}\lambda_{i}f_{i}\otimes e_{i}$, with $(e_{i})_{i=1}^{n}$ and $(f_{i})_{i=1}^{n}$ orthonormal sequences in $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$ respectively. Since $f_{i}\in\operatorname{ran}{Y}\subset\overline{\operatorname{ran}{B}}$, we can find $f_{i}^{m}\in\overline{\operatorname{ran}{B}}$ such that $\lVert f_{i}-f_{i}^{m}\rVert<1/n$ whenever $m\geq n$ and $n\in\mathbb{N}$. Let us define $Y^{(n)}\coloneqq\sum_{i=1}^{r}\lambda_{i}f_{i}^{n}\otimes e_{i}$. Then,
$\lVert Y-Y^{(n)}\rVert_{HS}^{2}=\sum_{i=1}^{r}\lambda_{i}\lVert f_{i}^{n}-f_{i}\rVert^{2}\leq r\lambda_{1}n^{-2}\rightarrow 0,\quad n\rightarrow\infty.$
As in the proof of Lemma 3.7 item (i), we see that $\hat{X}^{(n)}\coloneqq B^{\dagger}Y^{(n)}C^{\dagger}$ is a well-defined element of $\mathcal{X}_{r}$ satisfying (12) and $B\hat{X}^{(n)}C=Y^{(n)}$. In particular, $B\hat{X}^{(n)}C\rightarrow Y$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Using Lemma 3.6 twice and Proposition A.7, it follows that for $X\in\mathcal{X}_{r}$,
$\inf\{\lVert M-BXC\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}\}$ $\leq\lim_{n}\lVert M-B\hat{X}^{(n)}C\rVert_{HS}^{2}$
$=\lim_{n}\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\hat{X}^{(n)}C\rVert_{HS}^{2}+c$
$=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}^{2}+c$
$\leq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+c$
$=\lVert M-BXC\rVert_{HS}^{2}.$
By taking the infimum over $X\in\mathcal{X}_{r}$ we see that $\lVert M-B\hat{X}^{(n)}C\rVert_{HS}$ converges to $\inf\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{X}_{r}\}$.
3.3 Optimal approximation error
The quality of the optimal approximation (11) is addressed next.
Proposition 3.11.
Assume 3.1 and $\mathcal{Y}_{r}\not=\emptyset$. For any $\hat{X}$ given by (11) it holds that
$\lVert M-B\hat{X}C\rVert_{HS}^{2}=\lVert M\rVert_{HS}^{2}-\Delta,$
with $\Delta\coloneqq\sum_{i\leq r}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}$. We have,
$\Delta=\sum_{i\leq r}\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MC^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}})=\sum_{i\leq r}\sigma_{i}(C^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}}MC^{\dagger}C)$
If additionally $\operatorname{ran}{M}\subset\operatorname{dom}{B^{\dagger}}$, then we may write $\Delta=\sum_{i\leq r}\lambda_{i}(B^{\dagger}MC^{\dagger}CM^{*}B),$ where $\lambda_{i}(\cdot)$ denotes taking the $i$-th largest eigenvalue.
Proof.
By Lemma 3.7 item (ii), $B\hat{X}C$ is a rank-$r$ reduced SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. So, by Lemma 3.6,
$\lVert M-BXC\rVert_{HS}^{2}$ $=\lVert M\rVert_{HS}^{2}+\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\hat{X}C\rVert_{HS}^{2}-\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\rVert_{HS}^{2}$
$=\lVert M\rVert_{HS}^{2}+\sum_{i>r}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}-\sum_{i}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}$
$=\lVert M\rVert_{HS}^{2}-\Delta.$
By Proposition A.8, and the fact that $P_{\overline{\operatorname{ran}{B}}}$ and $P_{\ker{C}^{\perp}}$ are orthogonal projectors,
$\Delta=\sum_{i\leq r}\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})^{*})=\sum_{i\leq r}\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}M^{*}P_{\overline{\operatorname{ran}{B}}}),$
and similarly
$\Delta=\sum_{i\leq r}\sigma_{i}((P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})^{*}P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})=\sum_{i\leq r}\sigma_{i}(P_{\ker{C}^{\perp}}M^{*}P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}).$
Now we can use $P_{\ker{C}^{\perp}}=C^{\dagger}C$, c.f. (36). Finally, the assumption $\operatorname{ran}{M}\subset\operatorname{dom}{B^{\dagger}}$ and (37) imply $P_{\overline{\operatorname{ran}{B}}}M=BB^{\dagger}M$. Hence,
$T\coloneqq P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}M^{*}P_{\overline{\operatorname{ran}{B}}}=BB^{\dagger}MC^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}},$
and $T\in L_{2}(\mathcal{H}_{2})$, $T=T^{*}$, $T\geq 0$. In particular, the singular values of $T$ equal the eigenvalues of $T$. To compute $\Delta$ under the assumption $\operatorname{ran}{M}\subset\operatorname{dom}{B}^{\dagger}$, we may thus use the eigenvalues of $T$. Suppose that there exists nonzero $v$ such that $Tv=\lambda v$, so $v\in\operatorname{ran}{B}\subset\operatorname{dom}{B^{\dagger}}$. With $w\coloneqq B^{\dagger}v$ it holds that $w\not=0$ and by (35) and (37), $B^{\dagger}MC^{\dagger}CM^{*}Bw=B^{\dagger}Tv=\lambda w$. So $\lambda$ is an eigenvalue of $B^{\dagger}MC^{\dagger}CM^{*}B$. Conversely, suppose there exists nonzero $w$ such that $B^{\dagger}MC^{\dagger}CM^{*}Bw=\lambda w$, so $w\not\in\ker{B}$. With $v\coloneqq Bw$ it holds that $v\not=0$ and by (34) and (37), $Tv=BB^{\dagger}MC^{\dagger}CM^{*}Bw=B\lambda w=\lambda v$. So $\lambda$ is an eigenvalue of $T$. We conclude that
$\Delta=\sum_{i\leq r}\sigma_{i}(T)=\sum_{i\leq r}\lambda_{i}(T)=\sum_{i\leq r}\lambda_{i}(B^{\dagger}MC^{\dagger}CM^{*}B).$
4 Unbounded solution
We consider an example that shows that for a given approximation problem, all solutions $\hat{X}$ of (10) of the form (11) — and in fact all solutions of (10) — are unbounded. This example also shows it is possible that some solutions of the form (11) are bounded, while others are unbounded.
Let $\mathcal{H}$ be a separable Hilbert space and take $\mathcal{H}_{i}=\mathcal{H}$. Let $(e_{n})_{n}$ be an orthonormal basis of $\mathcal{H}$. Let $\0(\gamma)_{n},(\mu_{n})_{n}\in\mathbb{C}\backslash\{0\}$ converge to zero. We assume $(\mu_{n})_{n}$ is nonincreasing. Let $(\alpha_{n})_{n}\in\mathbb{C}$ such that $\lvert\alpha_{n}\rvert\rightarrow\infty$ and $(\alpha_{n}\gamma_{n})_{n}$ is square summable. For example, $\gamma_{n}=n^{-2}$ and $\alpha_{n}=n$ satisfy these conditions.
We may then define $w\coloneqq\sum_{n\geq 2}\alpha_{n}\gamma_{n}e_{n}\in\mathcal{H}$. Put $f_{1}\coloneqq w\lVert w\rVert^{-1}$ and $f_{2}\coloneqq e_{1}$ and extend $\{f_{1},f_{2}\}$ to an ONB $\{f_{n}\}_{n}$ of $\mathcal{H}$. We then set, for $h\in\mathcal{H}$,
$Mh=\sum_{n}\mu_{n}\langle h,f_{n}\rangle f_{n},$
$Ch=\sum_{n}\gamma_{n}\langle h,e_{n}\rangle e_{n}.$
Then $C$ is injective and $\operatorname{ran}{C}=\{h\in\mathcal{H}:\ \sum_{n}\gamma_{n}^{2}\langle h,e_{n}\rangle^{2}<\infty\}$. As $\operatorname{ran}{C}$ contains $\{e_{n}\}_{n}$, it is dense in $\mathcal{H}$, thus $\operatorname{ran}{C}^{\perp}=\{0\}$. Therefore, $\operatorname{dom}{C^{\dagger}}=\operatorname{ran}{C}$ and the Moore–Penrose inverse of $C$ is given, as in Example A.6, to be
$C^{\dagger}h=\sum_{n}\gamma_{n}^{-1}\langle h,e_{n}\rangle e_{n},\quad h\in\operatorname{dom}{C^{\dagger}}.$
Let us take $B=I$ for simplicity, so $P_{\overline{\operatorname{ran}{B}}}=I$. As $C$ is injective, $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}=M$. Since $(\mu_{n})_{n}$ is nonincreasing, for any $r\geq 1$, one of the rank-$r$ truncated SVDs of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ is then given by,
$(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}h=\sum_{n=1}^{r}\mu_{n}\langle h,f_{n}\rangle f_{n}.$ (18)
With this SVD for given $r\geq 1$, we now compute for $h\in\operatorname{dom}{C}^{\dagger}$,
$B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger}h=\sum_{n=1}^{r}\sum_{m=1}^{\infty}\mu_{n}\gamma_{m}^{-1}\langle h,e_{m}\rangle\langle e_{m},f_{n}\rangle f_{n}.$
As $f_{1}=w\lVert w\rVert^{-1}$, it follows for $h=e_{m}\in\operatorname{dom}{C^{\dagger}}$, $m\geq 2$,
$\lVert B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger}e_{m}\rVert^{2}$ $=\sum_{n=1}^{r}\big{\lvert}\mu_{n}\gamma^{-1}_{m}\langle e_{m},f_{n}\rangle\big{\rvert}^{2}$
$\geq\big{\lvert}\mu_{1}\gamma^{-1}_{m}\langle e_{m},f_{1}\rangle\big{\rvert}=\lvert\mu_{1}\rvert\lvert\alpha_{m}\rvert\lVert w\rVert^{-1}\rightarrow\infty$
as $m\rightarrow\infty$, showing that $B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger}$ is unbounded. This shows that (11) can indeed give unbounded solutions. If $\mu_{1}>\mu_{2}$ and $r=1$, then (11) gives a unique solution, which shows that (11) can also give only unbounded solutions. By Corollary 3.10 it follows that in this case in fact all solutions of (10) are unbounded.
Now let $r=1$ and $\mu_{1}=\mu_{2}$, so that $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ has at least two distinct rank-$1$ truncated SVDs. The rank-$1$ truncated SVD (18), i.e. $(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{1}=\mu_{1}\langle\cdot,f_{1}\rangle f_{1}$, yields the following solution, which is of the form (11),
$\hat{X}_{a}h\coloneqq\mu_{1}\langle C^{\dagger}h,f_{1}\rangle f_{1}=\mu_{1}\sum_{n}\gamma_{n}^{-1}\langle h,e_{n}\rangle\langle e_{n},f_{1}\rangle f_{1}=\mu_{1}\lVert w\rVert^{-1}\sum_{n\geq 2}\alpha_{n}\langle h,e_{n}\rangle f_{1},$
for $h\in\operatorname{dom}{C^{\dagger}}$. By again taking $h=e_{m}$, $\hat{X}_{a}$ is seen to be unbounded. As $f_{2}=e_{1}$, the rank-$1$ truncated SVD $\mu_{2}\langle\cdot,e_{1}\rangle e_{1}$ yields the solution
$\hat{X}_{b}h\coloneqq\mu_{2}\langle C^{\dagger}h,e_{1}\rangle e_{1}=\mu_{2}\gamma_{1}^{-1}\langle h,e_{1}\rangle e_{1},\quad h\in\operatorname{dom}{C^{\dagger}},$
which is bounded.
We have thus constructed two solutions of the form (11), yet one of these is bounded, in particular it has finite rank and has finite Hilbert–Schmidt norm, and one is unbounded.
5 Bounded solutions and approximations
Theorem 3.2 prescribes a set of solutions that are of the form (11). As shown in Section 4, this set can contain unbounded solutions. It is even possible that some solution of the form (11) are bounded and others are unbounded. It is natural to ask when all solutions are bounded. In Section 5.1 we give necessary and sufficient conditions for (11) to have only bounded solutions. In the case they are not bounded, we use the concept of outer inverses to construct suitable bounded approximations in Section 5.2.
5.1 Necessary and sufficient conditions for bounded solutions
We recall that Proposition A.5 item (iv) gives a sufficient condition for (11) to consist of only bounded solutions, namely that $\operatorname{ran}{C}$ is closed in $\mathcal{H}_{2}$. An even stronger condition is given next.
Lemma 5.1.
Let $C\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$. If $\ker{C}^{\perp}$ is finite-dimensional, then $\operatorname{ran}{C}$ is closed.
Proof.
Since $\overline{\operatorname{ran}{C^{*}}}=\ker{C}^{\perp}$ by Proposition A.3, it follows that $\operatorname{ran}{C^{*}}$ is finite-dimensional when $\ker{C}^{\perp}$ is finite-dimensional. But finite-dimensional spaces are closed. By Proposition A.4, then also $\operatorname{ran}{C}$ is closed. ∎
It follows that all $\hat{X}$ in (11) are bounded when $\mathcal{H}_{1}$ is finite-dimensional. We now discuss a range condition on $C$ that is not just sufficient, but also necessary for all $\hat{X}$ to be bounded.
Proposition 5.2.
In 3.1 suppose $\mathcal{Y}_{r}\not=\emptyset$. Let $\hat{X}$ be given by (11) and define $Z\coloneqq B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$, so $\hat{X}=ZC^{\dagger}$. The following are equivalent:
  1. (i)
    $\hat{X}$ is bounded,
  2. (ii)
    $\ker{\hat{X}}$ is closed in $\operatorname{dom}{C^{\dagger}}$,
  3. (iii)
    $C(\ker{Z}\cap\ker{C}^{\perp})$ is closed in $\operatorname{ran}{C}$,
Proof.
The equivalence of items (i) and (ii) is proven similarly as it is proven for linear functionals, i.e. when $r=1$, that continuity is equivalent to having a closed kernel. We outline a proof in the appendix as Lemma B.1. To show the equivalence between items (ii) and (iii) we compute $\ker{\hat{X}}$. As $\operatorname{ran}{C^{\dagger}}=\ker{C}^{\perp}$, c.f. Proposition A.5 item (i),
$\ker{\hat{X}}$ =$=\{k\in\operatorname{dom}{C^{\dagger}}:\ C^{\dagger}k\in\ker{Z}\}$
=$=\{k\in\operatorname{dom}{C^{\dagger}}:\ C^{\dagger}k\in\ker{Z}\cap\ker{C}^{\perp}\}$
=$=\{k_{0}+k_{\perp}\in\operatorname{ran}{C}\oplus\operatorname{ran}{C}^{\perp}:\ C^{\dagger}k_{0}\in\ker{Z}\cap\ker{C}^{\perp}\}.$
If $k_{0}\in\operatorname{ran}{C}$ is such that $C^{\dagger}k_{0}\in\ker{Z}\cap\ker{C}^{\perp}$, then $k_{0}=P_{\overline{\operatorname{ran}{C}}}k_{0}=CC^{\dagger}k_{0}\in C(\ker{Z}\cap\ker{C}^{\perp})$ by (37). Conversely, if $k_{0}=Cz$ for some $z\in\ker{Z}\cap{\ker{C}}^{\perp}$, then $z=P_{\ker{C}^{\perp}}z=C^{\dagger}Cz=C^{\dagger}k_{0}$ by (36). Thus,
$\ker{\hat{X}}=C(\ker{Z}\cap\ker{C}^{\perp})\oplus\operatorname{ran}{C}^{\perp}=U\oplus\operatorname{ran}{C}^{\perp},$
with $U\coloneqq C(\ker{C}\cap\ker{C}^{\perp})$. If $U\oplus\operatorname{ran}{C}^{\perp}\subset\operatorname{ran}{C}\oplus\operatorname{ran}{C}^{\perp}$ is closed, then $U\subset\operatorname{ran}{C}$ is closed. Since $\operatorname{ran}{C}^{\perp}$ is closed, the converse holds as well. This shows that items (ii) and (iii) are equivalent. ∎
We give a final sufficient condition, a lower bound condition, that is weaker than $C$ having closed range. Since $C$ is bounded, $\ker{C}$ is closed. We recall that the quotient space $\mathcal{H}_{1}/\ker{C}$ then is a Banach space with norm $\lVert h+\ker{C}\rVert=\inf\{\lVert h+c\rVert:\ c\in\ker{C}\}$. The quotient map $Q\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{1}/\ker{C})$, $h\mapsto h+\ker{C}$ is continuous. See for example [16, Section III.4].
Corollary 5.3.
Suppose that there exists $c>0$ such that $\lVert Ch\rVert\geq c\lVert h+\ker{C}\rVert$ for all $h\in\ker{Z}\cap\ker{C}^{\perp}$. Then $\hat{X}$ is bounded.
Proof.
Let us denote $W\coloneqq\ker{Z}\cap\ker{C}^{\perp}$. We show that $C(W)$ is closed in $\mathcal{H}_{4}$. Then $C(W)$ is also closed in $\operatorname{ran}{C}$ and we may apply Proposition 5.2. Now, we can decompose $C=\tilde{C}\circ Q$ for some $\tilde{C}\in\mathcal{B}(\mathcal{H}_{1}/\ker{C},\mathcal{H}_{2})$. Let $(w_{n})_{n}$ be a sequence in $W$ such that $(Cw_{n})_{n}=(\tilde{C}(w_{n}+\ker{C}))_{n}$ converges to $k\in\mathcal{H}_{4}$. We need to show $k=Cw$ for some $w\in W$. By assumption, $\lVert w_{n}-w_{m}+\ker{C}\rVert\leq c^{-1}\lVert C(w_{n}-w_{m})\rVert$, so $(w_{n}+\ker{C})_{n}$ is a Cauchy sequence in $\mathcal{H}_{1}/\ker{C}$. But $\mathcal{H}_{1}/\ker{C}$ is complete, so $w_{n}+\ker{C}$ converges to some $h+\ker{C}\in\mathcal{H}_{1}/\ker{C}$. As $W$ and $\ker{C}$ are orthogonal, their direct sum $W+\ker{C}$ is closed, c.f. [4, Section 9.2.(A)]. Hence $h+\ker{C}\in W+\ker{C}$, i.e. $h+\ker{C}=w+\ker{C}$ for some $w\in W$. Now, $Cw=Ch=\tilde{C}(h+\ker{C})=\lim_{n}\tilde{C}(w_{n}+\ker{C})=k$, by continuity of $\tilde{C}$. ∎
In case $C$ has closed range, the operator $\tilde{C}$ in the proof of Corollary 5.3 also has closed range. But $\tilde{C}$ is also injective, so it has a bounded inverse on its range, by the inverse mapping theorem, c.f. [16, Theorem III.12.5]. Hence
$\lVert h+\ker{C}\rVert=\lVert\tilde{C}^{-1}\tilde{C}h\rVert\leq\lVert\tilde{C}^{-1}\rVert\lVert Ch\rVert,\quad\text{for all }h\in\mathcal{H}_{1}.$ (19)
In particular, we conclude that the lower bound condition of Corollary 5.3 is satisfied with $c\coloneqq\lVert\tilde{C}^{-1}\rVert^{-1}$ in the case where $C$ has closed range. The lower bound condition in Corollary 5.3 need not hold everywhere on $\mathcal{H}_{1}$, so it is weaker than (19). Hence, the condition in Corollary 5.3 is strictly weaker than the condition that $\operatorname{ran}{C}$ is closed.
5.2 Bounded approximations
Even when the equivalent conditions of Proposition 5.2 do not hold, we can enforce bounded solutions by instead considering the problem
$\inf\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\},$ (20)
with the minimality property
$P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}=X.$ (21)
As $\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$, this infimum is bounded from below by the minimum in (10). In the case that $\operatorname{ran}{C}$ is closed, the problems (10)-(12) and (20)-(21) coincide by Remark 3.5. Let us therefore assume that $\operatorname{ran}{C}$ is not closed, so $\ker{C}^{\perp}$ is infinite-dimensional by Lemma 5.1. We now show that in this case the infimum (20) is not attained in general, but is still equal to the minimum in (10). To do so, we construct a minimizing sequence $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ to (20)-(21) such that $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$. We do this under the assumption that $\mathcal{Y}_{r}\not=\emptyset$ and hence that a solution $\hat{X}$ of the form (11) actually exists.
Given $\hat{X}$ of the form (11), we construct the approximating sequence $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ based on the approximation of the unbounded Moore–Penrose inverse $C^{\dagger}$ by outer inverses as proposed by [25]. We briefly repeat the construction and its properties here, but refer to [25] for elaboration.
Let $(Y_{n})_{n}$ be an increasing sequence of finite-dimensional subspaces of $\overline{\operatorname{ran}{C}}$ such that $\overline{\cup_{n}Y_{n}}=\overline{\operatorname{ran}{C}}$. We define $X_{n}\coloneqq C^{*}Y_{n}$ and let $P_{n}$, $Q_{n}$ be the orthogonal projectors onto $Y_{n}$ and $X_{n}$ respectively. We then set $C_{n}\coloneqq P_{n}C$, so $C_{n}$ has finite rank. It can be shown that $(X_{n})_{n}$ is an increasing sequence of finite-dimensional subspaces, that $\overline{\cup_{n}X_{n}}=\ker{C}^{\perp}$ and that $\ker{C_{n}}=X_{n}$, $\operatorname{ran}{C_{n}}=Y_{n}$ so that ${\left.\kern-1.2ptC_{n}\vphantom{\big{|}}\right|_{X_{n}}}:X_{n}\rightarrow Y_{n}$ is a boundedly invertible operator. Hence, we can define the finite-rank operator $C_{n}^{\#}:\mathcal{H}_{2}\rightarrow\mathcal{H}_{1}$ as,
$C_{n}^{\#}(y)\coloneqq\begin{cases}({\left.\kern-1.2ptC_{n}\vphantom{\big{|}}\right|_{X_{n}}})^{-1}(y),\quad y\in Y_{n},\\ 0,\quad y\in Y_{n}^{\perp}.\end{cases}$
The following result states the approximation property of $C^{\dagger}$ by $C_{n}^{\#}$, see [25, Theorem 2.1] for a proof.
Theorem 5.4.
The sequence $(C_{n})_{n}\subset\mathcal{B}_{00}(\mathcal{H}_{2},\mathcal{H}_{1})$ satisfies,
  1. (i)
    $\operatorname{dom}{C^{\dagger}}=\{h\in\mathcal{H}_{2}:\lim_{n}C_{n}^{\#}h\text{ exists}\},$
  2. (ii)
    $C^{\dagger}h=\lim_{n}C_{n}^{\#}h$ for $h\in\operatorname{dom}{C^{\dagger}}$,
  3. (iii)
    $Q_{n}C^{\dagger}h=C_{n}^{\#}h$ for $h\in\operatorname{dom}{C^{\dagger}}$,
  4. (iv)
    $C_{n}^{\#}CC_{n}^{\#}=C_{n}^{\#}$, i.e. $C_{n}^{\#}$ is an outer inverse of $C.$
Let us now, under the condition $\mathcal{Y}_{r}\not=\emptyset$, set
$\hat{X}_{n}\coloneqq B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C_{n}^{\#}.$ (22)
The next result directly implies that the infimum in (20) and minimum in (10) indeed agree.
Proposition 5.5.
In 3.1 suppose $\mathcal{Y}_{r}\not=\emptyset$ and let $\hat{X}$ be given by (11). With $\hat{X}_{n}$ defined in (22), it holds that $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$, $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$ and $B\hat{X}_{n}C\rightarrow B\hat{X}C$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Furthermore, $\hat{X}_{n}$ satisfies (21).
In Section 4 it is shown that it may happen that all solutions $\hat{X}$ of (10) are unbounded. Since the infimum in (20) and the minimum in (10) are equal by Proposition 5.5, it then follows that the infimum in (20) is not attained in this case. Indeed, if it were attained by some bounded $\tilde{X}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$, then $\tilde{X}$ would be a bounded minimiser of (10), contradicting the fact that in this case all solutions are unbounded.
Proof.
Let us first notice that, as in the proof of Lemma 3.7 item (i), we can use the first equation of (14) and the line preceding it to find that
$B\hat{X}C=(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}P_{\ker{C}^{\perp}},$
and we can use Lemma A.1 to find that $B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ is bounded. Now, being a composition of bounded operators, $\hat{X}_{n}$ is bounded. Also, $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$ since $C_{n}^{\#}\rightarrow C^{\dagger}$ pointwise on $\operatorname{dom}{C^{\dagger}}=\operatorname{dom}{\hat{X}}$ by Theorem 5.4.
As the $(X_{n})_{n}$ are increasing, we can take an ONB $(e_{i})_{i}$ of $\ker{C}^{\perp}$ such that for all $n$, $\{e_{1},\ldots,e_{n}\}$ is an ONB of $X_{n}$ and $e_{i}\in X_{n}^{\perp}$, $i>n$. Now, $Q_{n}P_{\ker{C}^{\perp}}=Q_{n}$ since $\ker{Q_{n}}^{\perp}=X_{n}\subset\overline{\operatorname{ran}{C}}$. By Theorem 5.4, $C_{n}^{\#}=Q_{n}C^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$, so with (36), $C_{n}^{\#}C=Q_{n}C^{\dagger}C=Q_{n}$, and using also (34),(36),(37) it follows that,
$B\hat{X}_{n}C$ $=BB^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C_{n}^{\#}C=(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}Q_{n},$
$P_{\ker{B}^{\perp}}\hat{X}_{n}P_{\overline{\operatorname{ran}{C}}}$ $=B^{\dagger}B\hat{X}_{n}CC^{\dagger}=B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}Q_{n}C^{\dagger}=\hat{X}_{n}\quad\text{on }\operatorname{dom}{C^{\dagger}}.$
The second equation shows that (21) holds, as $\operatorname{dom}{C^{\dagger}}\subset\mathcal{H}_{2}$ densely and $\hat{X}_{n}$ is continuous, while with the first we compute,
$\lVert B\hat{X}C-B\hat{X}_{n}C\rVert_{HS}^{2}$ $=\lVert(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}(P_{\ker{C}^{\perp}}-Q_{n})\rVert_{HS}^{2}$
$=\sum_{i>n}\lVert(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}e_{i}\rVert^{2}\rightarrow 0,$
as the tail of a sum that converges because $(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ is finite-rank and thus Hilbert–Schmidt. ∎
One could also start with an increasing sequence $(X_{n})_{n}$ and define $Y_{n}=CX_{n}$ instead of choosing $(Y_{n})_{n}$ and setting $X_{n}=C^{*}Y_{n}$. So $\operatorname{ran}{P_{n}}=C(\operatorname{ran}{Q}_{n})$ and we could try to construct the bounded approximations $\tilde{C}_{n}\coloneqq C^{\dagger}P_{n}$ to $C^{\dagger}$. If $\overline{\cup_{n}X_{n}}=\ker{C}^{\perp}$, then still $\overline{\cup_{n}Y_{n}}=\overline{\operatorname{ran}{C}}$ and the pointwise convergence $\tilde{C}_{n}\rightarrow C^{\dagger}$ holds on $\operatorname{dom}{C^{\dagger}}$. But $C^{\dagger}P_{n}Ce_{i}$ is not equal to 0 for $i>n$, so $\tilde{C}_{n}C=C^{\dagger}P_{n}C$ is no longer equal to $Q_{n}$, which is what we used to show the $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ convergence in Proposition 5.5. In fact, the sequence,
$\sum_{i>n}\lVert(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}(I-C^{\dagger}P_{n}C)e_{i}\rVert^{2}$
need not be converging in $n$. So the statement on $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ convergence will not hold in general for this alternative approximation procedure.
Remark 5.6 (Bounded approximations of unbounded approximate minimisers).
The approximate minimisers $\hat{X}^{(m)}$ constructed in Section 3.2 may be unbounded. The procedure outlined before Proposition 5.5 can be used for each $m$ to approximate $\hat{X}^{(m)}$ with bounded approximations $\hat{X}^{(m)}_{n}$. Then $B\hat{X}^{(m)}_{n}C\rightarrow B\hat{X}^{(m)}C$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ as $n\rightarrow\infty.$ In particular $B\hat{X}^{(m)}_{m}C\rightarrow Y$ for some rank-$r$ reduced SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. Thus $\hat{X}^{(m)}_{m}$ is a sequence of bounded approximate minimisers.
Remark 5.7 (Adjoint formulation).
Proposition 5.5 also allows us to pass to an adjoint formulation of (10). Let us write $\mathcal{X}_{r}(B,C)$ and $\mathcal{Y}_{r}(B,C)$ for $\mathcal{X}_{r}$ and $\mathcal{Y}_{r}$ defined in 3.1 to stress the dependence on $B$ and $C$. The adjoint problem then is
$\min\{\lVert M^{*}-C^{*}XB^{*}\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}(C^{*},B^{*})\},$ (23)
with the minimality property
$P_{\overline{\operatorname{ran}{C}}}XP_{\ker{B}^{\perp}}=X.$ (24)
We now discuss why one might want to solve (23) instead of (10). The same reasoning holds for solving (23)-(24) instead of (10)-(12). Note that under the assumption $\mathcal{Y}_{r}(B,C)\not=\emptyset$ and $\mathcal{Y}_{r}(C^{*},B^{*})\not=\emptyset$, we have by Proposition 5.5,
$\min\{\lVert M-BXC\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}(B,C)\}$ $=\inf\{\lVert M-BXC\rVert_{HS}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}$
$=\inf\{\lVert M^{*}-C^{*}X^{*}B^{*}\rVert_{HS}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}$
$=\inf\{\lVert M^{*}-C^{*}XB^{*}\rVert_{HS}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})\}$
$=\min\{\lVert M^{*}-C^{*}XB^{*}\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}(C^{*},B^{*})\}.$
Here we used that $(BXC)^{*}=C^{*}X^{*}B^{*}$ holds true when $B,X,C$ are bounded. When $\mathcal{Y}_{r}(B,C)=\emptyset$, the minimum of the left hand side becomes an infimum that is not attained. It is, however, still possible that $\mathcal{Y}_{r}(C^{*},B^{*})\not=\emptyset$. In this case the adjoint problem (23) still has a solution and allows us to recover the minimal Hilbert–Schmidt approximation error, even though (10) does not admit a minimal solution. Another reason for why (23) can be useful is that unlike (10), which has only unbounded solutions, (23) admits a bounded solution. In this case, the adjoint problem allows to recover the minimal Hilbert–Schmidt approximation error using a bounded operator. Section 6 shows an example of the latter case.
6 Application in Signal Processing, Reduced Rank Regression and Linear Operator Learning
In this section we describe how Theorem 3.2 can be leveraged to obtain results in infinite-dimensional settings. As example applications, we consider the fields of signal processing, reduced rank regression, and linear operator learning. We first recall some examples of approximation problems that arise in these applications, before describing a problem that includes these examples as special cases.
In the field of signal processing, one common problem is to reconstruct a random input signal $x\in\mathbb{R}^{m}$ from a random observation signal $y\in\mathbb{R}^{m}$ that is related to $x$. One can assume that each observation satisfies $y=Gx+\varepsilon$ for some known linear map $G$ and additive noise $\varepsilon$, or weaker assumptions can be made on the relationship between $x$ and $y$. In [51], one considers the problem of reconstructing the signal $x$ as $Ay$, assuming additional structure only in the form of a known covariance matrix of $(x,y)$. To do so, the reconstruction quality is assessed in the expected squared reconstruction error, and thus the objective is to minimise $\mathbb{E}\lVert x-Ay\rVert^{2}$ over $A$. As the observation signal can be noisy, the rank of $A$ is restricted in order to filter out this noise. For more generality and motivated by applications, one can also allow $x$, $y$, and the reconstruction $Ay$ to be weighted by matrices $W_{x},W_{y}$ and $W_{A}$. The problem then becomes to find the rank- constrained minimiser $A$ of $\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}$. In [54], one considers an extension of this problem in which not just one signal but sets of signals $\{x(\alpha)\in\mathbb{R}^{m},\ \alpha\in K\}$ and $\{y(\alpha)\in\mathbb{R}^{m},\ \alpha\in K\}$ are considered, where $K$ is a suitable index set that may be countably or uncountably infinite, e.g. $K=\mathbb{N}$ or $K=[0,1]$, and $x(\alpha)$ and $y(\alpha)$ are finite-dimensional random vectors in $\mathbb{R}^{m}$ for every $\alpha$. Below, we consider the settings in [51, 54] in the case of infinite-dimensional signals, that is, in case the signals $x(\alpha)$ and $y(\alpha)$ are function-valued for every $\alpha\in K$.
[…]
A closely related problem that arises in the setting of linear operator learning of [36] is to minimise $\mathbb{E}\lVert x-Ay\rVert^{2}$, where $x$ and $y$ are allowed to take values in a separable Hilbert space, and where the search space for $A$ is taken to be either the space of bounded operators or Hilbert–Schmidt operators between these Hilbert spaces. This has applications in various kinds of regression problems, such as functional regression with functional response; see [36]. In this section, we study the operator learning problem in the case that $A$ is bounded and the rank of $A$ is restricted and finite.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a countably generated probability space, so that $L^{2}(\Omega)$ is separable, c.f. [14, Proposition 3.4.5]. Let $F,G$ be real separable Hilbert spaces. We let $x\in L^{2}(\Omega;F)$ and $y\in L^{2}(\Omega;G)$ be random variables. Here, $L^{2}(\Omega;F)$ and $L^{2}(\Omega;G)$ denote Lebesgue-Bochner spaces, so integrability is in the sense of square Bochner integrability, see [26, Section 1.2]. These spaces are then separable. This follows for example from [26, Proposition 1.2.29] or because $L^{2}(\Omega;F)\simeq L^{2}(\Omega)\otimes F$, which is separable as the tensor product of separable Hilbert spaces, c.f [29, Assumption 2.1.(g)].
Problem 6.1.
Given weighting operators $W_{x}\in\mathcal{B}(F,\mathcal{H}_{1}),W_{A}\in\mathcal{B}(\mathcal{H}_{2},\mathcal{H}_{1})$ and $W_{y}\in\mathcal{B}(G,\mathcal{H}_{3})$ for real separable Hilbert spaces $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$, solve
$\min\{\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2},\ A\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})\}.$ (25)
Note that in infinite dimensions it is not a priori clear that the minimum in (10) is indeed attained for a bounded operator, c.f. Section 4, so 6.1 might not have a solution.
If ${F}=G=\mathbb{R}^{m}$, then we are back in the setting of [51]. If ${F}=G=L^{2}(K;\mathbb{R}^{m})$ with $(K,\mathcal{K},\lambda)$ a countably generated measure space with $\sigma$-finite measure $\lambda$, which is separable, then we are in the setting of [54], by interpreting the infinite set of signals $\{x(\alpha):\ \alpha\in K\}$ as one signal $x\in L^{2}(\Omega\times K;\mathbb{R}^{m})\simeq L^{2}(\Omega;L^{2}(K,\mathbb{R}^{m}))$. If $W_{x},W_{y}$ and $W_{A}$ are taken to be the identity, we are in the linear operator learning setting of [36] but with an additional rank restriction imposed on $A$. If $W_{x}=\Gamma=W_{A}$, $W_{y}=I$ and $x$ and $y$ are finite-dimensional and have mean 0, (25) reduces to the reduced rank setting of [27].
For $z\in L^{2}(\Omega;E_{1}),w\in L^{2}(\Omega;E_{2})$ we define the ‘uncentered covariance’ $C_{wz}\in\mathcal{B}(E_{1},E_{2})$ by
$\langle C_{wz}e,f\rangle\coloneqq\int_{\Omega}\langle w(\omega),f\rangle\langle e,z(\omega)\rangle\operatorname{d}\!{\mathbb{P}}(\omega),\quad e\in E_{1},f\in E_{2}.$
We also define $C_{z}\coloneqq C_{zz}$. Hence with $x$ and $y$ as in 6.1, $C_{x}$ and $C_{y}$ are positive and self-adjoint, while $C_{xy}=C_{yx}^{*}$. We collect some further facts on covariances in Lemma 6.2, which is proven in Appendix B.
Lemma 6.2.
Let $z\in L^{2}(\Omega;E_{1})$ and $w\in L^{2}(\Omega;E_{2})$. Then,
  1. (i)
    If $E_{1}=E_{2}$, then $\operatorname{tr}\,\left({C_{wz}}\right)=\mathbb{E}\langle w,z\rangle<\infty$ ,
  2. (ii)
    $C_{Qw,z}=QC_{w,z}$ for $Q\in\mathcal{B}(E_{2},E_{3})$ and any separable Hilbert space $E_{3}$,
  3. (iii)
    $C_{zw}=C_{w}^{1/2}UC_{z}^{1/2}$ for some $U\in\mathcal{B}(E_{1},E_{2})$ satisfying $\lVert U\rVert\leq 1$ and $U=P_{\overline{\operatorname{ran}{C_{w}^{1/2}}}}UP_{\overline{\operatorname{ran}{C_{z}^{1/2}}}}$.
Since $x,y$ in 6.1 are square integrable, we see from Lemma 6.2 item (i) that $C_{x}$, $C_{y}$ and $C_{xy}$ are all trace class. Thus $C_{x}^{1/2},C_{y}^{1/2}$ are Hilbert–Schmidt. We can now solve 6.1, assuming that $\operatorname{ran}{W_{A}}$ is closed. We note that in order to prove this result, we apply the adjoint formulation approach outlined in Remark 5.7.
Proposition 6.3.
Suppose that $W_{A}$ has closed range. Then a solution to 6.1 exists if and only if there exists a truncated SVD $(P_{\overline{\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\operatorname{ran}{W_{A}}})_{r}$ satisfying
$\operatorname{ran}{(P_{\overline{\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\operatorname{ran}{W_{A}}})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}.$ (26)
In that case, a solution is given by
$\hat{A}=W_{A}^{\dagger}\left((C_{y}^{1/2}W_{y}^{*})^{\dagger}(P_{\overline{\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\operatorname{ran}{W_{A}}})_{r}\right)^{*},$ (27)
which also satisfies the minimality property
$P_{\ker{W_{A}}^{\perp}}\hat{A}P_{\ker{C_{y}^{1/2}W_{y}^{*}}^{\perp}}=\hat{A}.$ (28)
Additionally, (27) uniquely specifies a solution of 6.1 that also satisfies (28) if and only if the truncated SVD in (27) is unique.
Proof.
Let $A\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})$. Let us also put $B\coloneqq C_{y}^{1/2}W_{y}^{*}\in\mathcal{B}(\mathcal{H}_{3},G),$ $C\coloneqq W_{A}^{*}\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2}),$ $N\coloneqq BA^{*}C$ and $M\coloneqq(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}$. Then $\operatorname{rank}\left({N}\right)\leq\operatorname{rank}\left({A^{*}}\right)=\operatorname{rank}\left({A}\right)\leq r$, so $N$ is Hilbert–Schmidt. By Lemma 6.2 item (iii), $C_{yx}=C_{y}^{1/2}UC_{x}^{1/2}$ with $U\in\mathcal{B}(F,G)$ and $\operatorname{ran}{U}\subset\overline{\operatorname{ran}{C}_{y}^{1/2}}$. Hence, with (36) and Proposition A.3,
$(C_{y}^{1/2})^{\dagger}C_{yx}=P_{\ker{C_{y}^{1/2}}^{\perp}}UC_{x}^{1/2}=P_{\overline{\operatorname{ran}{C_{y}^{1/2}}}}UC_{x}^{1/2}=UC_{x}^{1/2},$ (29)
so that $M$ is a well-defined Hilbert–Schmidt operator, as the product of the Hilbert–Schmidt operator $C_{x}^{1/2}$ and bounded operators. We can now use Lemma 6.2 items (i) and (ii) to compute
$\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}$ $=\operatorname{tr}\,\left({W_{A}AW_{y}C_{y}W_{y}^{*}A^{*}W_{A}^{*}}\right)+\operatorname{tr}\,\left({W_{x}C_{x}W_{x}^{*}}\right)$
$-\operatorname{tr}\,\left({W_{A}AW_{y}C_{yx}W_{x}^{*}}\right)-\operatorname{tr}\,\left({W_{x}C_{xy}W_{y}^{*}A^{*}W_{A}^{*}}\right).$
Multiplying (29) by $C_{y}^{1/2}$ from the left and using (34), we see that $C_{y}^{1/2}(C_{y}^{1/2})^{\dagger}C_{yx}=C_{yx}.$ As all Hilbert spaces are over $\mathbb{R}$, it holds that $\operatorname{tr}\,\left({S^{*}T}\right)=\operatorname{tr}\,\left({ST^{*}}\right)$ for $S,T\in L_{2}$. Therefore,
$\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}$ $=\operatorname{tr}\,\left({W_{A}AW_{y}C_{y}W_{y}^{*}A^{*}W_{A}^{*}}\right)$
$+\operatorname{tr}\,\left({W_{x}C_{x}W_{x}^{*}}\right)-2\operatorname{tr}\,\left({W_{A}AW_{y}C_{yx}W_{x}^{*}}\right)$
$=\lVert C_{y}^{1/2}W_{y}^{*}A^{*}W_{A}^{*}\rVert_{HS}^{2}+\lVert W_{x}C_{x}^{1/2}\rVert_{HS}^{2}$
$-2\operatorname{tr}\,\left({W_{A}AW_{y}C_{y}^{1/2}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}}\right)$
$=\lVert W_{x}C_{x}^{1/2}\rVert_{HS}^{2}+\lVert N\rVert_{HS}^{2}-2\operatorname{tr}\,\left({N^{*}M}\right)$
$=c+\lVert M-N\rVert_{HS}^{2}$
$=c+\lVert M-BA^{*}C\rVert_{HS}^{2}$ (30)
where $c\coloneqq\lVert W_{x}C_{x}^{1/2}\rVert_{HS}^{2}-\lVert M\rVert_{HS}^{2}$ does not depend on $A$. Since taking the adjoint is a bijective operation from $\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})$ onto $\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ and since $A^{**}=A$, we see from (30) that $A$ solves 6.1 if and only if $A^{*}$ solves
$\min\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}.$ (31)
Problem (31) does not have a solution when, with $\mathcal{X}_{r}$ as in Theorem 3.2, the problem
$\min\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{X}_{r}\}.$ (32)
does not have a solution, since $B_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$. If (32) does have a solution, then it has a solution $\hat{X}$ of the form (11) by Theorem 3.2, which then also solves (31). Indeed, by Proposition A.4, $\operatorname{ran}{W_{A}^{*}}$ is closed if and only if $\operatorname{ran}{W_{A}}$ is closed. By the assumption that $\operatorname{ran}{W_{A}}$ is closed and Proposition A.5 item (iv) it then follows that $(W_{A}^{*})^{\dagger}$ and hence $\hat{X}$ are bounded. Thus, 6.1, (31) and (32) are all equivalent, and $X$ solves 6.1 if and only if $X^{*}$ solves (31) and (32).
Using Theorem 3.2 we find that an optimal solution to (32) exists if and only if
$\operatorname{ran}{(P_{\overline{\operatorname{ran}{B}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\ker{C}^{\perp}})_{r}}\subset\operatorname{dom}{(C_{y}^{1/2}W_{y}^{*})^{\dagger}},$
for some truncated rank-$r$ SVD, which by Remark 3.4 is equivalent to (26). In that case, a solution $\hat{X}$ to (32) and (31) that also satisfies
$P_{\ker{B}^{\perp}}\hat{X}P_{\overline{\operatorname{ran}{C}}}=\hat{X},$
is given by
$\hat{X}=(C_{y}^{1/2}W_{y}^{*})^{\dagger}(P_{\overline{\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\ker{W_{A}^{*}}^{\perp}})_{r}(W_{A}^{*})^{\dagger}.$
This solution is unique if and only if the rank-$r$ truncated SVD in its definition is unique. Since $\hat{X}^{*}$ solves 6.1, the proof is completed after taking the adjoint and applying Proposition A.5 item (v) and Proposition A.3 to $W_{A}^{*}$. ∎
Passing to the adjoint in the formulation (32) allows us to use Theorem 3.2, as in this case Theorem 3.2 gives bounded solutions, which are precisely the ones we search for in 6.1. If we would not pass to the adjoint formulation and instead try to use Theorem 3.2 on
$\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}=c+\lVert M^{*}-C^{*}AB^{*}\rVert_{HS}^{2}$
then since $B^{*}$ does not have closed range, we would in general find an unbounded minimiser $\tilde{A}$. Since 6.1 considers bounded solutions, such an unbounded minimiser cannot be a solution. While such a minimiser $\tilde{A}$ might satisfy $\lVert M^{*}-C^{*}\tilde{A}B^{*}\rVert_{HS}<\lVert M-B\hat{A}C\rVert_{HS}$, the mean square error $\mathbb{E}\lVert W_{x}x-W_{A}\tilde{A}W_{y}y\rVert^{2}$ now is ill-defined in general. Indeed, it requires $W_{y}y\in\operatorname{dom}{\tilde{A}}$ almost surely, which is a restrictive condition and is not satisfied in general. In particular, if $W_{y}=I$ and $y$ has a nondegenerate Gaussian distribution, then $\operatorname{dom}{\tilde{A}}=\operatorname{ran}{C_{y}^{1/2}}$ is the Cameron–Martin space of $y$ which has zero measure in infinite dimensions, see e.g. [5, Theorem 2.4.7]. Thus, $W_{y}y\not\in\operatorname{dom}{\tilde{A}}$ almost surely.
When no solution to 6.1 exists, it can still be approximated using the technique described in Section 3.2 and Remark 5.6, applied to the adjoint formulation (32).
We can now study the problem equivalent to the linear operator learning setting, but with a rank restriction.
Corollary 6.4.
A solution to the problem
$\min\{\mathbb{E}\lVert x-Ay\rVert^{2},\ A\in\mathcal{B}_{00,r}(G,F)\}$ (33)
exists if and only if $(\operatorname{ran}{(C_{y}^{1/2})^{\dagger}C_{yx}})_{r}\subset\operatorname{ran}{C_{y}^{1/2}}$ for some rank-$r$ restricted SVD $(\operatorname{ran}{(C_{y}^{1/2})^{\dagger}C_{yx}})_{r}$. In that case, a solution that also satisfies the minimality property $\hat{A}P_{\ker{C_{y}}^{\perp}}=\hat{A}$ is given by
$\hat{A}=\left((C_{y}^{1/2})^{\dagger}(P_{\overline{\operatorname{ran}{C_{y}^{1/2}}}}(C_{y}^{1/2})^{\dagger}C_{yx})_{r}\right)^{*}.$
This solution is the only solution to (33) which satisfies $\hat{A}P_{\ker{C_{y}}^{\perp}}=\hat{A}$ if and only if the rank-$r$ reduced SVD is unique.
Proof.
We can apply Proposition 6.3 for the case $\mathcal{H}_{1}=\mathcal{H}_{2}=F$, $\mathcal{H}_{3}=G$ and $W_{A},W_{x},W_{y}$ equal to the identity operators. By (29) it holds that $\operatorname{ran}{(C_{y}^{1/2})^{\dagger}C_{yx}}\subset\overline{\operatorname{ran}{C_{y}^{1/2}}}$. The condition (26) therefore becomes $\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$. Finally, (28) becomes $AP_{\ker{C_{y}^{1/2}}^{\perp}}=A$. It remains to show that $\ker{C_{y}^{1/2}}=\ker{C_{y}}$. If $0=Ah=A^{1/2}A^{1/2}h$, then $A^{1/2}h\in\ker{A^{1/2}}=\operatorname{ran}{A^{1/2}}^{\perp}$ using Proposition A.3. As also $A^{1/2}h\in\operatorname{ran}{A^{1/2}}$ it follows that $\ker{A}\subset\ker{A^{1/2}}$. The reverse inclusion is immediate. ∎
With $U$ as in the proof of Proposition 6.3, the existence condition $\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$ is equivalent to $\operatorname{ran}{(UC_{x}^{1/2})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$.
Remark 6.5 (Condition comparison with [36]).
In [36] the analogous problem to (33) is studied, in which there is no rank restriction on $A$. There, it is shown in [36, Proposition 3.11] that a solution to this analogous problem exists if and only if the existence condition $\operatorname{ran}{C_{yx}}\subset\operatorname{ran}{C_{y}}$ holds, in which case the solution is given by $(C_{y}^{\dagger}C_{yx})^{*}$. Under the latter condition, for all $f\in F$ there exists $h\in G$ such that $C_{yx}f=C_{y}h$. Hence by (36) and Proposition A.3,
$(C_{y}^{1/2})^{\dagger}C_{yx}f=P_{\ker{C_{y}^{1/2}}^{\perp}}C_{y}^{1/2}h=P_{\overline{\operatorname{ran}{C_{y}^{1/2}}}}C_{y}^{1/2}h=C_{y}^{1/2}h\in\operatorname{ran}{C_{y}^{1/2}}.$
Hence,
$\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{(C_{y}^{1/2})^{\dagger}C_{yx}}\subset\operatorname{ran}{C_{y}^{1/2}}.$
We therefore see that the existence condition in [36] is stronger than the existence condition of Corollary 6.4. In particular, whenever one can find the optimal continuous linear operator $A$ without rank restriction, one can also find an optimal continuous and rank restricted operator. The converse does not hold in general.
Remark 6.6 (Relation between minimality condition and maximal kernel).
The minimality condition $\hat{A}=\hat{A}P_{\ker{C_{y}}^{\perp}}$ of Corollary 6.4 is equivalent to $\hat{A}(\ker{C}_{y})=0$. In [36, Proposition 3.11 and Remark 3.12(b)] it is shown that the solution of the unconstrained problem also satisfies this property, and in fact has maximal kernel among all solutions. The solution of the rank-constrained problem (33) can be seen to have maximal kernel among all solutions as well. Indeed, by applying Corollary 3.9 to (32) and taking the adjoint, we see that an arbitrary solution of (33) is of the form $\tilde{A}\coloneqq\hat{A}+T^{*}P_{\ker{C}_{y}}$ for some bounded $T:F\rightarrow G$. Now $\ker{T^{*}P_{\ker{C}_{y}}}^{\perp}\subset\ker{C}_{y}$. Hence $\ker{\tilde{A}}$ is maximal if and only if $T^{*}P_{\ker{C_{y}}}=0$, which holds if and only if $\tilde{A}=\hat{A}$. Furthermore, $\tilde{A}=\hat{A}$ holds if and only if $\hat{A}$ satisfies the minimality property $\hat{A}P_{\ker{C_{y}}^{\perp}}=\hat{A}$. Thus, the minimality property originating from (12) is exactly the maximal kernel property mentioned in [36, Proposition 3.11 and Remark 3.12(b)].
7 Conclusion
Generalised rank-constrained approximation in Frobenius norm of a matrix has had a myriad of applications, for example in signal processing, statistical inverse problems and data transmission systems. This finite-dimensional approximation problem has been analysed in [38, 20, 40], for example. In this work we have extended the analysis in the following ways.
In [20], a solution to the generalised rank-constrained approximation problem was proposed, which also satisfies a minimal norm property. We have shown that this property does not hold in general. We have proposed the modified minimality property (15) which is satisfied in general. The modified result in the finite-dimensional setting is given in Corollary 3.8.
Secondly, we have studied the infinite-dimensional version of the generalised approximation problem, studying existence, uniqueness and continuity of a solution. In particular, we have given necessary and sufficient conditions for existence and uniqueness of a solution. This is done in Theorem 3.2. Theorem 3.2 also gives explicit solutions that satisfy the minimality property. We have also shown that the proposed solutions — and in fact all solutions — may be discontinuous in Section 4, and have given necessary and sufficient conditions for the proposed solutions to be continuous in Section 5.1. In addition, we have constructed approximate solutions when no solution exists in Section 3.2, and continuous approximations to discontinuous solutions in Section 5.2.
Thirdly and finally, we have studied problems from signal processing, reduced rank regression and linear operator learning in infinite-dimensional settings. In particular, we have generalised the signal processing problem of [51, 54] in the case where both the unknown signal and the observation are finite-dimensional random vectors, to the case where both can be random functions. We have also studied the reduced rank regression setting of [28] to infinite dimensions but with Hilbert–Schmidt norm error instead of operator norm error. Finally, we have discussed the linear operator learning setting of [36], where in our case the operator to be learned is not only bounded or Hilbert–Schmidt but has finite rank. In these settings, we used Theorem 3.2 to find explicit solutions, and to give necessary and sufficient conditions for existence and uniqueness of solutions.
In summary, since finite-dimensional approximation problems often originate from a discretised version of an infinite-dimensional problem, our infinite-dimensional result paves the way for the analysis of various reduced rank approximation problems such as reduced rank linear operator learning in their native infinite-dimensional formulation.
8 Acknowledgements
The research of the authors has been partially funded by the Deutsche Forschungsgemeinschaft (DFG) Project-ID 318763901 – SFB1294.
Appendix A Elements from Functional Analysis
In this section we recall some key concepts and results and introduce notation. We consider specific properties of linear mappings from $\mathcal{H}$ to $\mathcal{K}$ for possibly infinite-dimensional separable Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ over $\mathbb{C}$ or $\mathbb{R}$. In particular, we recall the definition of certain operator classes in Section A.1, the Moore–Penrose inverse in Section A.2, the SVD theorem for compact operators and the Eckart–Young Theorem in Section A.3.
A.1 Linear Operators
To accommodate the treatment of unbounded operators, we introduce the following concept of a linear map, c.f also [16, Paragraph X.1].
A linear operator $T:\mathcal{H}\rightarrow\mathcal{K}$ is a function defined on a linear subspace of $\mathcal{H}$, denoted by $\operatorname{dom}{T}$, on which $T$ acts in a linear fashion. We also write $T:\operatorname{dom}{T}\subset\mathcal{H}\rightarrow\mathcal{K}$ to stress that $T$ is only defined on $\operatorname{dom}{T}$ and not necessarily on all of $\mathcal{H}$. When $\operatorname{dom}{T}\subset\mathcal{H}$ densely, we call $T$ ‘densely defined’. Without loss of generality, one can always assume that $T$ is densely defined when it is linear, as one can put $T$ equal to 0 on $\overline{\operatorname{dom}{T}}^{\perp}={\operatorname{dom}{T}}^{\perp}$ and $\operatorname{dom}{T}\oplus\operatorname{dom}{T}^{\perp}$ is dense in $\mathcal{H}=\overline{\operatorname{dom}{T}}\oplus\operatorname{dom}{T}^{\perp}$. Here, $U\oplus V$ denotes the direct sum of linear subspaces $U,V$ and $U^{\perp}$ the orthogonal complement of $U$. If also $S:\mathcal{H}\rightarrow\mathcal{K}$ is linear and $\operatorname{dom}{T}\subset\operatorname{dom}{S}$ and $S=T$ on $\operatorname{dom}{T}$, then we call $S$ an ‘extension’ of $T$. We define the kernel of $T$ by $\ker{T}\coloneqq\{h\in\operatorname{dom}{T}:\ Th=0\}$. If
$\lVert T\rVert\coloneqq\inf\{C>0:\ \lVert Th\rVert\leq C\lVert h\rVert,\ h\in\operatorname{dom}{T}\}<\infty$
then we say that $T$ is ‘bounded’, and call $\lVert T\rVert$ its ‘operator norm’. Otherwise, $T$ is called ‘unbounded’. As $T$ is linear, it is bounded if and only if it is continuous. If $T$ is bounded, then it has by continuity a unique bounded extension to $\overline{\operatorname{dom}{T}}$. If $T$ is bounded and $\dim\operatorname{ran}{T}<\infty$, then we say that $T$ is ‘finite-rank’ and define $\operatorname{rank}\left({T}\right)\coloneqq\dim\operatorname{ran}{T}$. If also $S:\mathcal{Z}\rightarrow\mathcal{H}$ is a linear operator for some separable Hilbert space $\mathcal{Z}$, then we define $T+S$ on $\operatorname{dom}{T+S}\coloneqq\operatorname{dom}{S}\cap\operatorname{dom}{T}$, and $TS$ on $\operatorname{dom}{TS}\coloneqq S^{-1}(\operatorname{dom}{T})$. If $S$ is finite-rank, then so is $TS$, which is collected below for future reference.
Lemma A.1.
Let $T:\mathcal{H}\rightarrow\mathcal{K}$ be linear. If $S:\mathcal{Z}\rightarrow\mathcal{H}$ is a finite-rank operator, then so is $TS$ and $\operatorname{rank}\left({TS}\right)\leq\operatorname{rank}\left({S}\right)$.
Proof.
As $\operatorname{rank}\left({S}\right)$ is finite, ${\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\operatorname{ran}{S}}}$ is bounded, since every linear operator is bounded on a finite-dimensional domain. Also,
$\operatorname{rank}\left({TS}\right)$ $=\operatorname{rank}\left({{\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\operatorname{ran}{S}}}}\right)=\dim{\operatorname{ran}{S}}-\dim{\ker{{\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\operatorname{ran}{S}}}}}$
$\leq\dim{\operatorname{ran}{S}}=\operatorname{rank}\left({S}\right),$
by the rank-nullity theorem in linear algebra, which can be applied because $\operatorname{dom}{{\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\operatorname{ran}{S}}}}=\operatorname{ran}{S}$ is finite-dimensional. ∎
The following is a classical example of an unbounded operator.
Example A.2.
Let $\mathcal{H}$ be an infinite-dimensional separable Hilbert space and $(e_{n})_{n}$ an orthonormal system in $\mathcal{H}$. Let $\0(\gamma_{n})_{n}\subset\mathbb{C}\backslash\{0\}$ be a sequence converging to zero. Let $\mathcal{D}\coloneqq\{h\in\mathcal{H}:\ \sum_{n}\gamma_{n}^{-2}\lvert\langle h,e_{n}\rangle\rvert^{2}<\infty\}$. We define the mapping,
$T:\mathcal{D}\subset\mathcal{H}\rightarrow\mathcal{H},\quad h\mapsto\sum_{n}\gamma_{n}^{-1}\langle h,e_{n}\rangle e_{n}.$
Then $T$ is linear, but not bounded, since $\lVert Te_{n}\rVert=\lvert\gamma_{n}\rvert^{-1}$ diverges as $n\rightarrow\infty$. Furthermore, $T$ is densely defined if and only if $(e_{n})_{n}$ forms an orthonormal basis. We can extend $T$ to a densely defined operator $\bar{T}$ on $\mathcal{H}$, by putting $\bar{T}h=0$ for $h\in\mathcal{D}^{\perp}$ and $\bar{T}h=Th$ for $h\in\mathcal{D}$. Then $\bar{T}$ is a well-defined linear operator on $\mathcal{D}\oplus\mathcal{D}^{\perp}$, which is dense in $\mathcal{H}$.
By $\mathcal{B}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H}$, $\mathcal{K})$, $\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$ and $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ we denote the sets of the linear operators $\mathcal{H}\rightarrow\mathcal{K}$ defined on all of $\mathcal{H}$ that are bounded, compact, finite-rank and have rank at most $r$, respectively. We endow these sets with the operator norm and note that all of them except for $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$ are linear spaces. We recall that an operator between Hilbert spaces is compact when it is a limit of finite-rank operators in operator norm. When $\mathcal{H}=\mathcal{K}$ we simply write $\mathcal{B}(\mathcal{H})$, $\mathcal{B}_{0}(\mathcal{H})$, $\mathcal{B}_{00}(\mathcal{H})$ and $\mathcal{B}_{00,r}(\mathcal{H})$. For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ we denote its Hilbert space adjoint by $T^{*}\in\mathcal{B}(\mathcal{K},\mathcal{H})$. That is, $T^{*}$ is the unique operator satisfying $\langle Th,k\rangle=\langle h,T^{*}k\rangle$ for $h\in\mathcal{H},k\in\mathcal{K}$. If $T=T^{*}$, then $T$ is ‘self-adjoint’. If $\langle Th,h\rangle\geq 0$ for all $h\in\mathcal{H}$, then $T$ is ‘positive’. If $T$ is positive and self-adjoint, there exists a unique positive self adjoint square root $T^{1/2}$ with $T^{1/2}T^{1/2}=T$. For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ arbitrary, we have that $T^{*}T$ is positive and self adjoint and can define $\lvert T\rvert\coloneqq(T^{*}T)^{1/2}$. We refer to [16, Sections II.2, II.4, VIII.3] for more on the above notions. The following relation between $T$ and $T^{*}$ holds. Note that it implies that $\ker{T}=\operatorname{ran}{T^{*}}^{\perp}$ and $\ker{T^{*}}=\operatorname{ran}{T}^{\perp}$.
Proposition A.3.
Let $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then $\overline{\operatorname{ran}{T}}=\ker{T^{*}}^{\perp}$ and $\overline{\operatorname{ran}{T^{*}}}=\ker{T}^{\perp}$.
This can be found in for example [16, Theorem VI.1.8]. We also recall part of the closed range theorem, c.f. [16, Theorem VI.1.10].
Proposition A.4.
Let $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then $\operatorname{ran}{T}$ is closed if and only if $\operatorname{ran}{T^{*}}$ is closed.
We also recall the definition of Hilbert–Schmidt operators and the Hilbert–Schmidt norm. If $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ satisfies $\sum_{i}\lVert Te_{i}\rVert^{2}<\infty$ for some orthonormal basis (ONB) $(e_{i})_{i}$ of $\mathcal{H}$, then we recall that $T$ is called a ‘Hilbert–Schmidt’ operator and $\lVert T\rVert_{HS}\coloneqq\left(\sum_{i}\lVert Te_{i}\rVert^{2}\right)^{1/2}$ is independent of the choice of $(e_{i})_{i}$. The linear space consisting of all Hilbert–Schmidt operators $L_{2}(\mathcal{H},\mathcal{K})\coloneqq\{T\in\mathcal{B}(\mathcal{H},\mathcal{K}):\ \sum_{i}\lVert Te_{i}\rVert^{2}<\infty\}$ becomes a Hilbert space, with inner product $\langle T,S\rangle_{HS}\coloneqq\sum_{i}\langle Te_{i},Se_{i}\rangle$, which is again independent of the choice of ONB. We call $\langle\cdot,\cdot\rangle_{HS}$ the ‘trace inner product’, and it induces the Hilbert–Schmidt norm $\lVert\cdot\rVert_{HS}$. For $h\in\mathcal{H},k\in\mathcal{K}$, we can interpret the tensor product $k\otimes h\in\mathcal{H}\otimes\mathcal{K}\simeq L_{2}(\mathcal{H},\mathcal{K})$ as the rank-1 map $\tilde{h}\mapsto\langle\tilde{h},h\rangle k$. If $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is the product of two Hilbert–Schmidt operators, then it is called ‘trace-class’. For a trace-class operator $T$ we define its trace $\operatorname{tr}\,\left({T}\right)\coloneqq\sum_{i}\langle Te_{i},f_{i}\rangle$, where $(e_{i})_{i}$ and $(f_{i})_{i}$ are any ONBs of $\mathcal{H}$ and $\mathcal{K}$ respectively. The value of $\operatorname{tr}\,\left({T}\right)$ does not depend on the choice of these ONBs. If $T\in L_{2}(\mathcal{H},\mathcal{K})$, then $\lvert T\rvert^{1/2}$ is trace-class. We recall that a trace-class operator is Hilbert–Schmidt, and a Hilbert–Schmidt operator is compact. Finally, we notice that by the above definitions, $\langle T,S\rangle_{HS}=\operatorname{tr}\,\left({S^{*}T}\right)$ for $S,T\in L_{2}(\mathcal{H},\mathcal{K})$. For a discussion on Hilbert–Schmidt and trace class operators, we refer to [15, Section 3.18].
A.2 Moore–Penrose inverse
We recall the definitions of the Moore–Penrose inverse on Hilbert spaces, see also for example [18, Section 2.1] or [24, Section 3.5].
Let $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. The Moore–Penrose inverse $T^{\dagger}$ of $T$ is given by a densely defined linear operator $\mathcal{K}\rightarrow\mathcal{H}$ constructed as follows. The continuous linear map $\tilde{T}\coloneqq{\left.\kern-1.2ptT\vphantom{\big{|}}\right|_{\ker{T}^{\perp}}}:\ker{T}^{\perp}\rightarrow\mathcal{K}$ is injective and hence invertible on $\operatorname{ran}{T}$. The inverse $\tilde{T}^{-1}$ is thus well-defined on $\operatorname{ran}{T}$. Also, $\tilde{T}^{-1}$ is bounded if and only if $\operatorname{ran}{T}$ is closed. Let $\operatorname{dom}{T^{\dagger}}\coloneqq\operatorname{ran}{T}\oplus\operatorname{ran}{T}^{\perp}$, which is dense in $\overline{\operatorname{ran}{T}}\oplus\operatorname{ran}{T}^{\perp}=\mathcal{K}$. Since every $x\in\mathcal{H}$ has a unique decomposition $x=h+k$ for $h\in\operatorname{ran}{T}$ and $k\in\operatorname{ran}{T}^{\perp}$, we can define the action of $T^{\dagger}$ on $\operatorname{dom}{T^{\dagger}}$ unambiguously via $T^{\dagger}(h+k)\coloneqq\tilde{T}^{-1}h$. So $T^{\dagger}:\mathcal{K}\rightarrow\mathcal{H}$ is linear.
The following holds, the proof of which can be found in [18, Propositions 2.3 and 2.4] for items (i), (ii), (iii) and (iv). Item (v) is a special case of [4, Theorem 9.3.2 (g)], as the conclusion also holds for closed operators and for operators that do not satisfy the closed range condition.
Proposition A.5.
Let $\mathcal{H},\mathcal{K}$ be Hilbert spaces and $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then,
  1. (i)
    $\operatorname{ran}{T^{\dagger}}=\ker{T}^{\perp}$,
  2. (ii)
    $T$ and $T^{\dagger}$ satisfy the ‘Moore–Penrose equations’:
    $TT^{\dagger}T$ $=T,$ (34)
    $T^{\dagger}TT^{\dagger}$ $=T^{\dagger},$ (35)
    $T^{\dagger}T$ $=P_{\ker{T}^{\perp}},$ (36)
    $TT^{\dagger}$ $={\left.\kern-1.2ptP_{\overline{\operatorname{ran}{T}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{T^{\dagger}}}},$ (37)
  3. (iii)
    $T^{\dagger}$ is a closed operator, i.e. it has closed graph,
  4. (iv)
    $T^{\dagger}$ is bounded if and only if $T$ has closed range,
  5. (v)
    If $T$ has closed range, then $(T^{*})^{\dagger}=(T^{\dagger})^{*}$.
We consider an example in which the Moore–Penrose operator is unbounded.
Example A.6.
Assume the setting of Example A.2 with $(e_{n})_{n}$ an orthonormal basis. Hence $\mathcal{D}\subset\mathcal{H}$ densely, $\mathcal{D}^{\perp}=\{0\}$ and $\bar{T}=T$. We define $S\in\mathcal{B}_{0}(\mathcal{H})$ as
$Sh=\sum_{n}\gamma_{n}\langle h,e_{n}\rangle e_{n},\quad h\in\mathcal{H}.$
As $\gamma_{n}\rightarrow 0$, $S$ is indeed compact as the limit of the finite-rank operators $h\mapsto\sum_{n=1}^{N}\gamma_{n}\langle h,e_{n}\rangle e_{n}$ in $B(\mathcal{H},\mathcal{K})$ when $N\rightarrow\infty$. Also, $\operatorname{ran}{S}=\mathcal{D}$ and $\ker{S}=\{0\}$. If $k\in\mathcal{D}$, then $h\coloneqq\sum_{n}\gamma_{n}^{-1}\langle k,e_{n}\rangle e_{n}=Tk\in\mathcal{H}$ and $k=Sh$. Thus $S^{\dagger}:\mathcal{D}=\mathcal{D}\oplus\mathcal{D}^{\perp}\rightarrow\mathcal{H}$ is given by ${T}$.
A.3 Singular Value Decomposition and Eckart–Young Theorem
Let $T\in\mathcal{B}_{0}(\mathcal{H},\mathcal{K})$. We define $\mathcal{I}\coloneqq\{1,\ldots,\dim{\operatorname{ran}{T}}\}$. Thus, if $T$ has infinite-dimensional range, then $\mathcal{I}$ contains infinitely many elements. If $T$ has finite-dimensional range, then $\mathcal{I}$ contains $\operatorname{rank}\left({T}\right)$ elements. Then there exists a nonincreasing sequence $(\sigma_{i})_{i\in\mathcal{I}}\subset[0,\infty)$, an orthonormal system $(e_{i})_{i\in\mathcal{I}}$ of $\mathcal{H}$ and an orthonormal system $(f_{i})_{i\in\mathcal{I}}$ of $\mathcal{K}$ such that
$T=\sum_{i\in\mathcal{I}}\sigma_{i}f_{i}\otimes e_{i},\quad\text{i.e. }Th=\sum_{i\in\mathcal{I}}\sigma_{i}\langle h,e_{i}\rangle f_{i},\quad h\in\mathcal{H},$
see e.g. [24, Section 4.3]. We call such a series expansion of $T$ an SVD of $T$. The $i$-th singular value $\sigma_{i}$ is also denoted by $\sigma_{i}(T)$ to express its dependence on $T$. The summation can be finite or countably infinite, depending on whether $T$ is finite-rank or not. For any $r\in\mathbb{N}$ we can truncate the sum
$(T)_{r}\coloneqq\sum_{i=1}^{r}\sigma_{i}f_{i}\otimes e_{i}.$ (38)
This truncation is uniquely defined if and only if one of the following conditions holds: $r\geq\dim{\operatorname{ran}{T}}$ or $\sigma_{r}>\sigma_{r+1}$. The first condition expresses the scenario in which $T$ is a finite-rank operator with rank at most $r$, so that $(T)_{r}=T$. The second condition, $\sigma_{r}>\sigma_{r+1}$, is a spectral gap condition, and implies that $T$ is not finite-rank or that $T$ is finite-rank but its rank is at least $r$. Both conditions hold simultaneously if and only if $\operatorname{rank}\left({T}\right)=r$.
The following result is known as the Eckart–Young Theorem in the case of finite-dimensional matrices. See for example, [24, Theorem 4.4.7].
Proposition A.7 (Eckart–Young Theorem).
Let $T\in L_{2}(\mathcal{H},\mathcal{K})$. Then $(T)_{r}$ solves the problem,
$\min\{\lVert T-S\rVert_{HS}:\ S\in\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})\}.$
The SVD of self-adjoint normal operators reduces to $T=\sum_{i\in\mathcal{I}}\lambda_{i}e_{i}\otimes e_{i}$ with real eigenvalues $\lambda_{i}\rightarrow 0$ and orthonormal basis $(e_{i})_{i}$. This follows from the spectral theorem for compact self-adjoint operators, c.f. [16, Theorem II.5.1]. The Hilbert–Schmidt norm can now be calculated in the following way.
Proposition A.8.
Let $T\in L_{2}(\mathcal{H},\mathcal{K})$, then $\lVert T\rVert_{HS}^{2}=\sum_{i}\sigma_{i}(T^{*}T)=\sum_{i}\sigma_{i}(TT^{*})$.
Proof.
We have $T^{*}T=\sum_{i}\sigma_{i}(T^{*}T)e_{i}\otimes e_{i}$, with ONB $(e_{i})_{i}$ of eigenvectors and real eigenvalues $(\sigma(T^{*}T)_{i})_{i}$. Hence, $\lVert T\rVert_{HS}^{2}=\operatorname{tr}\,\left({T^{*}T}\right)=\sum_{i.j}\langle T^{*}Te_{i},e_{j}\rangle=\sum_{i}\sigma_{i}(T^{*}T)$. The proof for $TT^{*}$ is analogous. ∎
Appendix B Proofs of results
Lemma B.1.
Let $\mathcal{H},\mathcal{K}$ be normed spaces and $T:\operatorname{dom}{T}\subset\mathcal{H}\rightarrow\mathcal{K}$. Suppose that $r\coloneqq\dim{\operatorname{ran}{T}}<\infty$. Then $T$ is continuous if and only if $\ker{T}$ is closed in $\operatorname{dom}{T}$.
Proof.
If $T$ is continuous, then $\ker{T}=T^{-1}(\{0\})$ is closed as $\{0\}$ is closed. For the converse, note that $\ker{T}\subset\operatorname{dom}{T}$ closed implies that $\operatorname{dom}{T}/\ker{T}$ is a normed space and the quotient map $Q:\operatorname{dom}{T}\rightarrow\operatorname{dom}{T}/\ker{T}$, $h\mapsto h+\ker{T}$ is continuous, see [16, Section III.4]. Let us define $S:\operatorname{dom}{T}/\ker{T}\rightarrow\mathcal{K}$, $S(h+\ker{T})=Th$ Then $S$ is a well-defined and linear map satisfying $T=S\circ Q$. So $\operatorname{ran}{S}=\operatorname{ran}{T}$ and if $S(h+\ker{T})=0$, then $Th=0$ so that $h\in\ker{T}$ and $S$ is injective. This implies that $\dim{\operatorname{dom}{T}/\ker{T}}=r$. So $S$ is continuous as linear map defined on a finite-dimensional space, c.f. Proposition [16, Propositions III.3.4]. We conclude that $T$ is continuous as a composition of continuous maps. ∎
See 6.2
Proof.
Let $(e_{i})_{i}$ be an ONB of $E_{1}$. Applying the Cauchy-Schwarz inequality twice, we get
$\mathbb{E}\sum_{i}\langle w,e_{i}\rangle\langle e_{i},z\rangle=\mathbb{E}\langle w,z\rangle\leq\mathbb{E}\lVert z\rVert\lVert w\rVert\leq(\mathbb{E}\lVert z\rVert^{2})^{1/2}(\mathbb{E}\lVert w\rVert^{2})^{1/2}<\infty.$
Hence by Fubini’s Theorem,
$\operatorname{tr}\,\left({C_{wz}}\right)=\sum_{i}\langle C_{wz}e_{i},e_{i}\rangle=\sum_{i}\mathbb{E}\langle w,e_{i}\rangle\langle e_{i},z\rangle=\mathbb{E}\sum_{i}\langle w,e_{i}\rangle\langle e_{i},z\rangle=\mathbb{E}\langle z,w\rangle.$
This proves item (i). For item (ii), we note that for all $e\in E_{1}$ and $f\in E_{2}$,
$\langle C_{Qw,z}e,f\rangle=\mathbb{E}\langle f,Qw\rangle\langle e,z\rangle=\mathbb{E}\langle Q^{*}f,w\rangle\langle e,z\rangle=\langle C_{wz}e,Q^{*}f\rangle=\langle QC_{wz}e,f\rangle.$
The statement of item (iii) is proven in [1, Theorem 1] in the case of centered covariances. However, as noted in [29, Proof of Theorem 5.4], the proof also works for uncentered covariances, showing that item (iii) holds. ∎